### v1

# SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures

Abstract.
Training large Deep Neural Network (DNN) models requires thousands of GPUs for days or weeks at a time. At these scales, failures are frequent and can have a big impact on training throughput. Restoring performance using spare GPU servers becomes increasingly expensive as models grow.
SlipStream is a system for efficient DNN training in the presence of failures, without using spare servers. It exploits the functional redundancy inherent in distributed training systems – servers hold the same model parameters across data-parallel groups – as well as the bubbles in the pipeline schedule within each data-parallel group. SlipStream dynamically re-routes the work of a failed server to its data-parallel peers, ensuring continuous training despite multiple failures. However, re-routing work leads to imbalances across pipeline stages that degrades training throughput. SlipStream introduces two optimizations that allow re-routed work to execute within bubbles of the original pipeline schedule. First, it decouples the backward pass computation into two phases. Second, it staggers the execution of the optimizer step across pipeline stages. Combined, these optimizations enable schedules that minimize or even eliminate training throughput degradation during failures. We describe a prototype for SlipStream and show that it achieves high training throughput under multiple failures, outperforming recent proposals for fault-tolerant training such as Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 ×, respectively.
copyright: none
1. Introduction
Deep Neural Networks (DNNs) are consistently achieving milestone results in domains such as natural language processing (OpenAI, 2022), speech recognition (Hinton et al., 2012), and computer vision (Krizhevsky et al., 2017). Since the rapid growth of DNNs is a major contributor to recent breakthroughs (Villalobos et al., 2022), we are now in a global race to develop increasingly-large foundation models (Bommasani et al., 2021; Hoffmann et al., 2022). Both proprietary (Brown et al., 2020; Achiam, 2023; Wang et al., 2023) and open-source (Touvron et al., 2023; Databricks, 2024) models have tens to hundreds of billions of parameters. Training such models requires uninterrupted access to thousands of accelerators (e.g., GPUs) for several weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020). DNN training is essentially a large supercomputing job that relies on hybrid parallelism, which concurrently leverages data, tensor, and pipeline parallelism strategies (Zheng et al., 2022).
As the scale and duration of training increases, so does the likelihood of encountering failures (Gupta et al., 2017; Jeon et al., 2019; Jiang et al., 2024; Eisenman et al., 2022; Weng et al., 2022; Zu et al., 2024). While scale-out workloads, such as distributed databases and analytics, are designed to work around server failures, this is not yet the case for DNN training. The rigid parallelization of DNN training implies that each failure can set off a domino effect. All resources are forced to idle while a failed node is replaced and the job is re-optimized and re-started on the new hardware configuration. For example, the training of OPT-175B included 178,000 GPU-hours of wasted time due to various malfunctions (Le Scao et al., 2022).
Fault tolerance in DNN training involves three issues: fault detection, checkpointing, and efficient execution in the presence of faults. Large-scale training systems implement comprehensive monitoring of the health of hardware and software components, so that faults or stragglers are quickly detected and diagnosed (Jiang et al., 2024). Recent research has focused on reducing the overhead of periodic checkpoints of the training job state (Eisenman et al., 2022; Mohan et al., 2021; Korthikanti et al., 2023). Efficient execution in the presence of faults has received less attention. A common approach in industry is to maintain a reserve of spare GPU servers that will replace failed servers when needed. While conceptually simple, this approach is expensive at scale. As the frequency of faults increases, so does the number of spares.
The alternative approach is to continue training with the subset of resources available. This is similar to how scale-out systems approach resilience to failures (Dean and Ghemawat, 2004). Data parallelism provides one simple implementation as it creates full replicas of the model across groups of GPUs. When one GPU server fails, we take all the servers in its data-parallel group offline and continue training the remaining model replicas. Unfortunately, this approach increases the blast radius of the failure and its impact to performance. When training a 530B GPT model using the hybrid-parallel scheme of Megatron-LM (Narayanan et al., 2021), a single node failure would force 280 GPUs to go offline and cause a roughly 11% drop in training throughput.
Two recent projects have exploited pipeline parallelism instead. Bamboo (Thorpe et al., 2023) relies on redundant computation and executes each pipeline stage on two nodes even in the fault-free case. When a node fails, the alternative nodes for its stages will do the work. Bamboo’s drawback is the drop in fault-free training throughput due to the redundant computations and added memory pressure on each GPU node. Oobleck (Jang et al., 2023) avoids overheads in fault-free execution by pre-computing a number of templates for pipeline parallelism that use a different number of nodes. As nodes fail, Oobleck replaces the original pipeline template with one that uses fewer nodes. The performance challenge for Oobleck is balancing work across heterogeneous pipelines as the pipelines with fewer nodes can become stragglers. Both Bamboo and Oobleck require significant re-configuration as nodes fail or rejoin, which can be a challenge as failure rates increase.
In this paper, we introduce SlipStream, a novel scheme for resilient training with hybrid-parallel systems. SlipStream enables efficient execution in the presence of failures, without the need for spares and without any impact on model accuracy. Similar to Bamboo and Oobleck, we exploit pipeline parallelism. But unlike these previous proposals that handle failures within a single pipeline, SlipStream utilizes the inherent functional redundancy across pipelines in hybrid-parallel training systems. SlipStream re-routes the micro-batches of failed nodes to peer nodes that process the same pipeline stage for the model in the other data-parallel groups. Peer nodes store the same model parameters and can handle the extra work without the need for parameter re-shuffling.
If done naively, SlipStream’s adaptive pipelining increases training time and memory usage due to the additional work for the peers of failed nodes. SlipStream uses a series of optimizations that exploit unique characteristics of hybrid-parallel training to eliminate all or most of the inefficiencies. SlipStream uses the bubbles in the pipeline schedule of peers to execute the additional work at low or no overhead. These bubbles typically appear during the start-up and cool-down phase of the pipeline schedule, but not during the steady-state when most additional work must occur. SlipStream overcomes this challenge by decoupling back propagation for micro-batches into two distinct gradient computations, one relative to the input and one relative to the parameters. The two steps are independently scheduled in order to better leverage bubbles in the the cool-down portion of the pipeline schedule. The selective application of decoupled back propagation exploits the imbalance in memory usage across stages in hybrid-parallel training while avoiding an increase in peak memory pressure. Finally, SlipStream staggers the execution of the optimizer tasks across pipeline stages in order to leverage bubbles from the start-up phase of the pipeline schedule for the next training iteration.
We implemented SlipStream on top of the DeepSpeed training framework (Rasley et al., 2020). The key SlipStream element is the Planner that uses dynamic programming (DP) and mixed integer linear programming (MILP) to find the best pipeline schedule in the presence of multiple failures. The Planner runs offline and calculates efficient schedules for up to a certain number of failures. These schedules are employed as needed during training. We evaluated SlipStream using DNNs with billions of parameters like GPT-3 (Brown et al., 2020), using both real-world experiments and simulations of large training systems. We show that SlipStream supports high throughput training even at high failure counts, e.g. $\geq$10of the overall system. SlipStream outperforms Oobleck by 1.46×1.46$\times$1.46 × in training scenarios with realistic GPU failures. SlipStream’s advantage stems from its efficient scheduling of re-routed work and its ability to avoid major re-shuffling of model parameters upon failures. SlipStream outperforms Bamboo by 1.64×1.64$\times$1.64 ×. It is also able to efficiently train much larger models than Bamboo, which requires large memory overheads for its fault tolerance mechanism.
2. Background and Motivation
2.1. Distributed DNN Training
Figure 1. Illustration of hybrid parallelism. Pipeline stages are denoted with different colors. Within each pipeline stage, operators are partitioned through tensor parallelism. The global batch is split into micro-batches across pipelines.
State-of-the-art Deep Neural Network (DNN) models consist of tens to hundreds of billions of parameters and are trained on datasets with many trillions of tokens (Brown et al., 2020; Achiam, 2023; Touvron et al., 2023; Le Scao et al., 2022; Zhang et al., 2022; Fedus et al., 2021; Lepikhin et al., 2020). Their training requires thousands of GPUs for days or weeks (Hoffmann et al., 2022; Smith et al., 2022; Naumov et al., 2020; Kaplan et al., 2020). For example, Llama-3 was trained on 15 trillion tokens, using two clusters of 24K GPUs (AI, 2024).
DNN training is parallelized using three primary forms of parallelism. Data parallelism (DP) processes subsets of the input data in parallel across GPU groups, each of which stores the entire model (Sergeev and Del Balso, 2018; Goyal et al., 2017; You et al., 2017). Data parallelism requires high network bandwidth for the all-reduce operations that reconcile model parameters across all replicas at the end of iteration steps. Tensor and pipeline parallelism are forms of model parallelism that facilitate training large models by sharding the model across GPUs. Tensor parallelism (TP) partitions the parameters of each layer across GPUs in order to parallelize each linear algebra operation within a layer (Shoeybi et al., 2019). It incurs high communication costs that are not easily hidden by computation due to frequent all-reduce operations in both the forward and backward pass. Pipeline parallelism (PP) divides the model into sequential groups of layers or stages. Micro-batches of data are processed in parallel in a pipelined manner across these stages (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023). Pipeline parallelism requires lower network bandwidth as stages only communicate activations and gradients at layer boundaries. However, it achieves lower compute utilization as the number of stages increases due to pipeline dependencies and bubbles (idle slots) in the pipeline schedule.
Large-scale training systems balance these tradeoffs by utilizing all three forms of parallelism as shown in Figure 1 (Narayanan et al., 2021; Li and Hoefler, 2021; Liu et al., 2023; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022). This is known as hybrid-parallelism. In the most common case, systems such as DeepSpeed (Rasley et al., 2020) or Alpa (Zheng et al., 2022) use a combination of tensor parallelism within a multi-GPU server, pipeline parallelism across multi-GPU servers, and data parallelism across pipelines. Hybrid parallelism enables large model training with graceful scaling and reasonable training times (weeks to few months) using optimized clusters with 1000s of densely-connected GPUs.
2.2. Fault Tolerance in Distributed DNN Training
Large training systems include thousands of GPUs, CPUs, memory chips, networking chips, cables of various types, and power conversion and cooling devices. Scale allows for high performance but also leads to frequent faults ranging from software errors to full hardware malfunctions. The Mean Time Between Failure (MTBF) for large training systems can be minutes (Gupta et al., 2017; Weng et al., 2022; Jiang et al., 2024; He et al., 2023). For example, large-scale training clusters at Microsoft see a failure every $\approx 45$ minutes (Jeon et al., 2019).
In contrast to scale-out frameworks like Map-Reduce (Dean and Ghemawat, 2004), which quickly adjust to dynamic changes in resource availability, DNN training operates like a supercomputing job. It relies on fixed sharding and parallelization strategies and gang-scheduled execution. All computational resources must concurrently run uninterrupted, making the system highly susceptible to disruptions from any failure. For example, Meta encountered over 100 hardware failures while training OPT-175B, resulting in the loss of 178,000 GPU-hours (Le Scao et al., 2022; Bekman, 2022). Similar failure rates have been reported by ByteDance (Jiang et al., 2024), Alibaba (He et al., 2023), LAION (Beaumont, 2022), Microsoft (Jeon et al., 2019) and Google (Zu et al., 2024). At this scale, faults are regular occurrences that require systematic optimizations to ensure resilient and efficient training (He et al., 2023; Jiang et al., 2024; Jeon et al., 2019; Weng et al., 2022).
Enhancing fault tolerance in distributed training involves three key issues: fast error detection, efficient checkpointing, and robust execution under faults. Fast error detection is crucial for minimizing downtime and preventing the spread of errors. Some errors can be quickly detected through encoding and redundancy mechanisms in the hardware (e.g., error correction codes) or software timeouts. Other errors are detected if they eventually lead to irregular spikes in the loss function optimized during training (Beaumont, 2022). The latter category includes silent data corruptions (Bacon, 2022; Dattatraya Dixit et al., 2022). Checkpointing, which saves the state of the model at regular intervals, allows for recovery from the recent stable state without restarting the entire training from scratch. Since naive checkpointing to remote storage leads to long pauses in the training process, there has been considerable research in improving its efficiency (Mohan et al., 2021; Eisenman et al., 2022; Wang et al., 2023; Jiang et al., 2024) as discussed in Section 6.
2.3. Efficient DNN Training in Presence of Failures
Our work focuses on robust training in the presence of faults. A common approach is to have a reserve of spares ready to replace servers with failed GPU servers. To reduce recovery time, spare servers can also be hot standbys with their memory loaded with model parameters at the end of each iteration. The key advantage of spare servers is that, upon recovery, training efficiency returns to its pre-failure state. However, this approach becomes increasingly costly as systems scale up, fault frequency increases, and more spares are needed. Spare servers must connect to the network of the training cluster with a similar bandwidth profile as the servers they will be called to replace, which incurs additional networking costs and typically requires that spares are allocated in coarse-grained groups (Jouppi et al., 2023). Hot standby servers should be provisioned according to the full capacity of a data parallel group, determined by multiplying pipeline parallelism ($PP$) by tensor parallelism ($TP$). For instance, training a 530B GPT model requires 280 spare GPUs on standby, which results in an 11% increase in training costs (Narayanan et al., 2021). Spare servers can potentially run small training jobs while on (non hot) standby to amortize cost. However, the job eviction latency needed to free these resources upon a failure can cause long stalls for large training jobs occupying thousands of GPUs.
The alternative approach is to continue training with the largest possible subset of functional resources. Ideally, when $x\%$ of GPU servers fail, training can continue at no less than $(100-x)\%$ of the fault-free throughput. The simplest method, sometimes referred to as elastic batching, is to drop a whole data parallel (DP) group when a node fails (Wang et al., 2018; Lin et al., 2019). The remaining data parallel groups can proceed with the same parallelization scheme. While straightforward, this method has substantial drawbacks. A single node failure results in $PP\times TP$ nodes going offline, decreasing the total throughput by $1/DP$ (e.g., 17% when training a 1000B parameter GPT model (Narayanan et al., 2021)). Moreover, this requires reducing the training batch size or recompiling the program to optimize each data parallel group for a larger portion of the batch. This can disrupt the balance and efficiency of the entire training job.
Figure 2. Adaptive pipelining when $W_{1\textunderscore 2}$ fails. The output from worker $W_{1\textunderscore 1}$ is redirected to workers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$.
(a) Fault-Free 1F1B Schedule
(b) Adaptive Schedule when $W_{1\textunderscore 2}$ fails.
Figure 3. Hybrid-parallel training with 3 data-parallel pipelines, 4 pipeline stages, and 6 micro-batches per iteration.  3(a) shows a 1F1B fault-free schedule.  3(b) shows how Adaptive Pipelining re-routes micro-batches failed workers to functional peers.
Two recent projects have instead utilized pipeline parallelism for efficient training in the presence of failures without spares. Bamboo (Thorpe et al., 2023) draws inspiration from RAID (Patterson et al., 1988) and designs a pipeline for redundant computation (RC). Each pipeline stage is redundantly computed on two nodes, even in the fault-free case. When a node fails, its backup node steps up and computes the forward and backward passes of the failed node. Bamboo uses RC even in the absence of faults. While RC tries to hide some overhead in pipeline bubbles, its extra computational overhead still significantly drops training throughput (see Section 5). RC also replicates the state of each pipeline stage in two nodes, increasing GPU memory pressure and limiting its scalability. Finally, Bamboo needs to restart with a full reconfiguration from a checkpoint for as few as two failures, e.g., when two adjacent nodes fail.
Oobleck (Jang et al., 2023) adapts pipeline parallelism for resilient execution but has no overhead in the fault-free case. It relies on pre-computed templates for pipeline parallelism. All templates are logically equivalent yet physically heterogeneous. Each template has a different number of stages, a different micro-batch configuration, and occupies a different number of nodes. A pipeline with failed node(s) is replaced with a new pipeline from a template that requires fewer nodes. However, templates with fewer nodes have lower training throughput. To avoid having the slowest pipeline slowdown the overall synchronous training job, Oobleck attempts to distribute the global mini-batch proportionally to compute power. This is not always possible and the straggling pipelines will inflate training time. Moreover, each pipeline re-configuration is in the critical path of training and can become an overhead if nodes fail frequently. Re-configuration is also needed when nodes re-join. Finally, Oobleck treats each heterogeneous pipeline as a black-box. It fails to observe that there are bubbles across all pipelines that can help hide some or all of the overheads due to node failures.
3. SlipStream Techniques
SlipStream aims to support efficient distributed training in the presence of faults. It requires no spare servers and has no impact on model accuracy compared to fault-free training. SlipStream can tolerate multiple hardware failures and maintains training throughput proportional to the number of functional servers available. SlipStream is optimized for fast recovery as failures do not require significant re-shuffling of model parameters between functional nodes.
This section reviews the key SlipStream techniques: Adaptive Pipelining, Decoupled BackProp, and a Staggered Optimizer. Section 4 presents the SlipStream system design.
3.1. Adaptive Pipelining: Working Around Failures
SlipStream exploits two key properties in hybrid-parallel training systems. First, there is functional redundancy across data-parallel pipelines. The GPUs that process the same pipeline stage hold identical parameters and only differ in the micro-batches they process. In Figure 2, for example, workers $W_{0\_2}$, $W_{1\_2}$, and $W_{2\_2}$ are peers that hold identical parameters for stage 2 of the 4-stage pipeline. Second, there are bubbles (idle slots) in the pipeline schedule for each worker. The 1F1B pipeline schedule (Narayanan et al., 2019) in Figure 3(a) has 9 bubbles in the repeating 27-slot schedule for worker $W_{0\_2}$.
Adaptive Pipelining exploits functional redundancy and bubbles by dynamically re-routing micro-batches from a failed worker to its functioning peers in other data-parallel pipelines. We aim to use the bubbles in peers’ schedules to process the micro-batches for the failed worker at low performance penalty. We evenly distribute micro-batches across all functional peers. Since all pipelines perform a synchronized all-reduce at the end of each iteration, load balancing ensures that no single worker (and thus pipeline) is overloaded, delaying the progress of the entire training iteration.
Figure 2 shows an example with three data parallel, 4-stage pipelines. Worker $W_{1\_2}$ fails. Upon detecting the failure, Adaptive Pipelining will redirect its input to workers $W_{0\_2}$ and $W_{2\_2}$. These two peer workers will process micro-batches received from worker $W_{1\_1}$ in addition to their regular pipeline load. The output for these additional micro-batches will be sent back to worker $W_{1\_3}$, the original recipient of the output of the failed worker $W_{1\_2}$. All other workers operate as in the fault-free schedule. In essence, Adaptive Pipelining repairs the functionality of pipeline 1 by exploiting bubbles in the data parallel peers of failed worker $W_{1\_2}$. Note that Adaptive Pipelining requires no model parameter re-shuffling or re-partitioning across workers in order to resume after a failure. Hence recovery is fast, unlike schemes like Oobleck that require re-configuring an entire pipeline upon a failure (Jang et al., 2023).
Figure 3 shows the detailed 1F1B schedule for the system in Figure 2. After the failure, the forward pass micro-batch 7 for failed worker $W_{1\_2}$ is re-assigned to its functional peer $W_{0\_2}$ at time step 3 ( 1). The subsequent output is forwarded back to worker $W_{1\_3}$ at time step 4 ( 2). Similarly, the forward pass of micro-batch 10 is re-routed to worker $W_{2\_2}$ at time step 12 ( 3). The backward pass operates similarly in the opposite direction. Worker $W_{2\_2}$ receives from worker $W_{1\_3}$ the gradients for micro-batch 12 at time step 22 ( 4). The output gradients are passed to worker $W_{1\_1}$ at time step 24 ( 5). The overall mathematical computation remains unchanged from the fault-free 1F1B schedule, ensuring that Adaptive Pipelining does not impact model convergence.
Figure 4. Forward and Backward pass for an operator.
The scheduling challenge: In theory, the execution schedule of the two peer workers has enough bubbles to do the computations for failed node $W_{1\_2}$ as there are 18 idle slots in total in Figure 3(a). However, these bubbles are concentrated in the warm-up and cool-down phase of the 1F1B schedule. Adaptive Pipelining needs to reroute micro-batches to peer workers $W_{0\_2}$ and $W_{2\_2}$ mostly in the middle of their steady-state schedule, which is optimized to be bubble-free for efficiency (Narayanan et al., 2019). As shown in Figure 3(b), this has a significant performance impact. The additional micro-batches to peer workers and the dependencies between forward and backward pass computations lead to 36 time steps per iteration. In other words, the failure of 8.3% of workers (1 out of 12) leads to a 33% slowdown in iteration time.
3.2. Decoupled BackProp: Filling Unused Bubbles
Figure 5. Optimized schedule with Decoupled BackProp when worker $W_{1\textunderscore 2}$ fails.
Decoupled BackProp addresses the scheduling problem of Adaptive Pipelining. It separates the backward pass into two distinct phases, allowing for more flexible scheduling of the extra load in the peers of a failed worker. Figure 4 shows that the backward pass calculates two distinct outputs: the gradients relative to the input ($B_{Input}$) and the gradients relative to parameters (weights) for that pipeline stage ($B_{Weight}$). Conventionally, $B_{Input}$ and $B_{Weight}$ are coupled as a unified calculation. This coupling lengthens dependencies between pipeline stages. The backward pass for stage $i$ must wait for the completion of both $B_{Input}$ and $B_{Weight}$ from stage $i+1$, despite stage $i$ requiring only $B_{Input}$ for its backward computations. $B_{Weight}$ can be deferred to improve the overall pipeline schedule. Thus, Decoupled BackProp splits the backward pass into two distinct tasks.
(a) Fault-free 1F1B Schedule
(b) Schedule for Adaptive Pipelining with Decoupled BackProp when worker $W_{1\textunderscore 2}$ fails.
(c) Schedule with the addition of the Staggered Optimizer.
Figure 6. The Staggered Optimizer allows SlipStream to optimize pipeline schedule across training iterations.
Figure 5 explains how Decoupled BackProp reduces the overheads introduced by Adaptive Pipelining. It shows the same example as in Figure 3 with the addition of Decoupled BackProp. The forward pass for the micro-batch 7 is unchanged ( 1 and 2). However, instead of waiting on a coupled backward pass (both $B_{Input}$ and $B_{Weight}$) from the previous stage to complete, which causes a ripple effect of dependencies throughout the pipeline, subsequent stages can continue processing new micro-batches using just the $B_{Input}$ from their predecessors. For example, worker $W_{0\_2}$ can run the backward pass for micro-batch 9 ( 3) and immediately pass the $B_{Input}$ to worker $W_{1\_3}$ ( 4) without computing $B_{Weight}$. $B_{Weight}$ is deferred to a later time ( 5) to reduce pipeline stalls, bringing the overall overhead down to just two time-steps (7.4% overhead with 8.3% failed workers).
Because $B_{Weight}$ is dependence-free, it can largely be deferred until the end of the training iteration. Hence, we can take advantage of idle slots in the cool-down phase of the 1F1B schedule for $B_{Weight}$ computations, freeing slots in the steady stage for the re-routed work of the failed worker. This enables SlipStream to absorb the additional computational demands imposed by failures with minimal overheads.
The memory challenge. Unfortunately, Decoupled BackProp increases memory pressure. Decoupling the computation of $B_{Weight}$ requires that intermediate data is stored for potentially extended periods on each worker. For instance, as shown in Figure 5, decoupling the backward pass for micro batch 1 necessitates retaining its intermediate data until time step 25. To avoid memory exhaustion, SlipStream applies Decoupled BackProp selectively only when it can mitigate the overheads of Adaptive Pipelining. We also capitalize on the observation that memory imbalance exists among the pipeline stages (Kim et al., 2023). To ensure that all pipeline stages are fully utilized, earlier stages need to allocate additional memory to process more forward micro-batches than later stages (Narayanan et al., 2019). By exploiting this available surplus memory, SlipStream can effectively offset some of the memory demands of Decoupled BackProp without incurring additional costs.
3.3. Staggered Optimizer: Accessing More Bubbles
Decoupled BackProp makes efficient use of the cool-down bubbles. However, as seen in Figure 6, the warm-up bubbles remain underutilized due to their placement after the synchronous optimizer step of the previous iteration. To exploit warm-up bubbles, we make the critical observation that optimizer steps for different pipeline stages are independent of each other. The Staggered Optimizer thus shifts the timing of the optimizer step across pipeline stages to better utilize the bubbles that occur during the warm-up phase of the subsequent training iteration. Effectively, this staggering provides SlipStream more bubbles in the cool-down phase to hide the overhead of compensating for failed workers.
Figure 6(c) shows how combining Staggered Optimizer with Adaptive Pipelining and Decoupled BackProp results in zero overhead over the fault-free 1F1B schedule in the running example. By staggering the optimizer step, later pipeline stages can move bubbles from the warm-up phase of the second iteration to the cool-down phase of the first iteration. Thus, peers that need to process additional micro-batches from failed workers within their steady-state schedule (i.e, workers $W_{0\_2}$ and $W_{2\_2}$) can further defer their weight gradient computations towards the cool-down phase. Workers for earlier stages (e.g., $W_{0\_0}$ for stage 0) can continue with the optimizer step, ensuring that the entire pipeline does not stall. Worker $W_{0\_0}$ can start iteration 2 at exactly the same time step as the fault-free 1F1B schedule, even with failures.
Put together, these three mechanisms allow SlipStream to flexibly optimize the use of computational resources across pipelines, ensuring that each stage operates near its capacity and maintains consistent utilization throughout training.
3.4. Supporting Multiple Failures
SlipStream’s techniques support multiple failures. SlipStream guarantees continued training up to $DP-1$ simultaneous worker failures, where $DP$ is the number of data-parallel pipelines, as one functional worker exists for each pipeline stage. SlipStream can probabilistically sustain more failures, as long as there is at least one functional worker per stage across all data parallel pipelines. Figure 7(b) shows how SlipStream can recover from more than $DP-1$ failures. Despite 8 simultaneous failures – 2/3 of the GPUs! – SlipStream continues training.
(a) The loss of an entire DP group requires checkpoint restoration.
(b) SlipStream can continue training if a functional worker exists per stage.
Figure 7. Example of SlipStream’s fault tolerance guarantees.
As workers are repaired, SlipStream can re-insert them back into the adapted pipeline schedule, reducing the probability of high counts of concurrent failures. In the rare event that simultaneous failures affect a single data parallel group as shown in Figure 7(a), SlipStream resorts to current resilience solutions: calculate an efficient hybrid-parallel scheme for the remaining nodes, configure it across all workers by restoring from a recent checkpoint, and resume training at the speed the new parallelism scheme allows.
Multi-GPU Servers and Tensor Parallelism. Training systems commonly employ servers with multiple GPUs. A favoured configuration involves HGX/DGX servers with 8 GPUs connected in an all-to-all manner using NVLink and NVSwitches (DGX, 2024). Similar to Oobleck and Bamboo, we consider a whole server as the unit of failure. This approach is pragmatic because servicing even one faulty GPU board necessitates taking the entire server offline. Moreover, software faults, CPU issues, connectivity problems, and power or cooling issues can also incapacitate a whole server. We also need to account for the level of tensor parallelism. Most training systems apply tensor parallelism within a multi-GPU server to leverage the high-bandwidth NVLink connections. In cases where tensor parallelism spans multiple servers, SlipStream, along with Oobleck and Bamboo, identifies the collective group of tensor-parallel servers as a unit of failure.
4. SlipStream Design
4.1. System Overview
SlipStream consists of three key components, the Profiler, the Executors, and critically the Planner.
Profiler. When a large training job is first submitted, SlipStream runs a short profiling job to collect key performance statistics such as the average micro-batch latency for the forward and backward passes, memory requirements for activations and gradients, and inter-node communication bandwidth. The profiling job executes a small number of training iterations, 100 by default, and typically takes a few minutes. These statistics are used by the Planner.
Planner. The Planner generates near-optimal schedules for the training job under various failure scenarios. It uses dynamic programming (DP) and mixed integer linear program (MILP) formulations that implement the SlipStream techniques (Section 3). We generate a pipeline schedule for each number of tolerated simultaneous failures. Each plan is agnostic to which specific worker(s) fail. It encodes the specific pipeline schedule including micro-batch assignment, the order of forward and backward tasks (e.g., ForwardPass, BackwardWeightPass), and the communication plan (e.g., RecvActiv, SendGrad, etc.). The plans are stored in distributed fault-tolerant storage (e.g., etcd) to be used by runtime Executors.
By default, we generate up to $DP-1$ plans, which accounts for the $DP-1$ simultaneous failures that SlipStream can always handle. Since SlipStream can probabilistically sustain many more failures, we can also generate plans for up to a user-defined fault tolerance threshold. The Planner runs relatively quickly, requiring 53 minutes to generate plans for up to 512 concurrent failures for a cluster of 2048 GPUs. Nevertheless, the training job can begin training in parallel as most plans will not be needed immediately.
Executors. SlipStream operates with an Executor on each GPU node, tasked with implementing the training plan specific to that node. Overseeing these Executors, a centralized Coordinator orchestrates the overall training operation by distributing the appropriate execution plans to each Executor. Additionally, the Coordinator actively monitors the status of all Executors to ensure their continuous operation. In the event of a failure, the Coordinator reassigns the tasks impacted by failure to the most suitable location, as advised by the Planner (refer to Section 4.2.1), and then directs the remaining Executors to resume training according to the updated plan that now reflects the updated count of non-operational workers. Training resumes from the iteration during which the failure was identified. If failures are not promptly detected, restoring a checkpoint from remote storage may become necessary. To enhance reliability, the Coordinator itself is safeguarded against failures through active replication.
4.2. SlipStream Planner
Given the cluster configuration, training job, profiling statistics, and set of failed workers, the Planner calculates a valid pipeline schedule that minimizes the training iteration latency using the fault-free workers. The Planner uses two phases, Normalization followed by Solution.
SlipStream generates an execution plan for a given number of failures, regardless of which worker(s) fails, as solving ILP problems for the combinatorial number of possible failure locations is intractable. The first phase normalizes the specific failures by migrating failed nodes to a pre-determined and ideal set of locations within the pipeline schedule. For example, if any single node fails in the example in Figure 7, the Planner will always swap the node with an ideal location, e.g. $W_{2\_3}$. The Planner generates a mapping of the ideal migration locations for each possible number of simultaneous failures. Unlike current solutions which require a complete reconfiguration of a data-parallel pipeline, normalization only requires a point-to-point copy of model parameters to swap the location of two workers in the pipeline for each failure. The Solution Phase leverages an MILP to derive an optimized schedule for each normalized case.
4.2.1. Normalization Phase
The intuition behind the Normalization Phase is to a) distribute failures across different peer groups (to maximize fault tolerance and load balance extra work across peer groups) and b) move failures to later pipeline stages with more bubbles (to maximize the opportunity to hide overheads).
Normalization relies on dynamic programming (DP) to determine the ideal pipeline location to migrate faults for given a number of failures. We use a DP algorithm as it can greedily determine the ideal migration strategy with very low runtime ($O(NF^{2})$ complexity for all solutions up to $F$ failures with $N$ stages per pipeline). The DP algorithm is based on a model that calculates the cost in terms of the extra bubbles used to accommodate re-routed micro-batches.
DP Formulation
Algorithm 1 Normalization Phase DP Algorithm
1:$N\leftarrow$ Number of pipeline stages.
2:$F\leftarrow$ Total number of failures.
3:$C\leftarrow$ an Nx(F+1) array $\triangleright$ Costs
4:$A\leftarrow$ an Nx(F+1) array $\triangleright$ Assignments
5:for $i\in\{0,\dots,N-1\}$ do
6:     for $f\in\{0,\dots,F\}$ do
7:         if $i==0$ then
8:              $C[i][f]=cost(i,f)$
9:              $A[i][f]=[f]$
10:         else
11:              $C[i][f]=\min_{x\leq f}(C[i-1][f-x]+cost(i,x))$
12:              $x=\operatorname*{arg\,min}_{x\leq f}(C[i-1][f-x]+cost(i,x))$
13:              $A[i][f]=concat(A[i-1][f-x],x)$
14:         end if
15:     end for
16:end for
17:$R=A[N-1][F]$
18:return $R$
Given the total number of failures to handle, $F$, and the total number of pipeline stages within a data parallel pipeline, $N$, the DP algorithm returns a list $R$ of length $N$, such that $\Sigma_{i}R[i]=F$. Each index $R[i]$ specifies the ideal number of failures that is assigned to pipeline stage $i$ across all data parallel pipelines; the specific pipeline assignment does not affect performance and can be arbitrary. For example, if $R[3]=2$ for the example in Figure 7(a), the ideal migration would be to swap the failed nodes with any two nodes in stage $3$, for example $W_{0\_3}$ and $W_{2\_3}$.
For each DP step, $C[i][f]$ represents the total cost, in terms of bubbles used, to handle $f$ failures at all stages between 00 and $i$. Meanwhile, each $A[i][f]$ contains a list that denotes the portion of the $f$ failures assigned to each stage from 00 to $i$, as described for $R$ – in other words $len(A[i][f])==i+1$. Each DP step uses a recurrence relation that aims to minimize the cost of handling $f$ failures within the first $i$ pipeline stages across all of the data parallel pipelines, while $A[i][f]$ denotes the assignment of failures that achieves this cost. The ideal assignment to handle $F$ failures across all workers is simply $R=A[N-1][F]$.
To estimate the cost, $cost(i,x)$ is a function that returns the extra bubbles used to handle the extra micro-batches from $x$ failures in the $i$-th stage (i.e., peer group). While we can use the MILP described in Section 4.2.2 to estimate the cost, the Planner can also rely on a heuristic to estimate the cost in order to avoid longer solve times (as Algorithm 1 requires $O(NF)$ calls to the cost function). The heuristic captures the benefit of the techniques described in Section 3. It creates a pipeline schedule that is similar to 1F1B, except that the heuristic’s 1F1B backward pass only schedules $B_{input}$. Then, when there exists a gap (both in terms of time and memory) that is able to fit $B_{weight}$ (or the $F$ and $B_{input}$ stages are exhausted), $B_{weight}$ operations are scheduled opportunistically.
4.2.2. Solution Phase
Next, the Planner uses the normalized failure locations and the profiled statistics as inputs to an MILP which performs schedule optimization. At a high level, the MILP determines how to schedule re-routes micro-batches across peers of a failed worker alongside regular micro-batches in those peers (Adaptive Pipelining). The formulation accounts for the latency for communicating gradients across workers, task dependencies in forward and backward passes, and the memory usage in each worker. The definition of task dependencies also allows the solution to leverage Decoupled BackProp and Staggered Optimizer techniques.
MILP Formulation
We aim to minimize the makespan of a single training iteration which is repeated throughout the training process, while the constraints are task dependencies and the memory capacity of workers.
Notation. An operation within a training iteration can be explicitly identified across workers using a 5-tuple $(i,j,k,c,k_{s})$. $i$ represents the pipeline stage within each data parallel pipeline, while $k$ represents the operation’s original data parallel pipeline (prior to failures). $j$ is the micro-batch ID in the training iteration. $c$ identifies the different operations for each micro-batch, $c\in\{F,B_{input},B_{weight}\}$. Finally, $k_{s}$ denotes the peer pipeline in the event that a micro-batch originally intended for $k$ runs on $k_{s}$ (note that $k$ can equal $k_{s}$). For example, a micro-batch originally intended to execute on $W_{2\_3}$, but was re-routed to a peer $W_{1\_3}$ in Figure 7, would have $i=3$, $k=2$, and $k_{s}=1$. Thus, $(i,k_{s})$ uniquely identifies the worker for a given operation.
Inputs. The MILP takes as input the micro-batch to worker assignments in an iteration. Specifically, the Planner takes the normalized failure locations and assigns all micro-batches corresponding to failed workers evenly across peer workers within the same peer group, alongside the original (non-peer) micro-batches. Thus, the Planner produces a binary mapping $S_{i,j,k}^{k_{s}}\in\{0,1\}$, which indicates if a micro-batch $(i,j,k)$ should be executed on pipeline $k_{s}$. Note that $S_{i,j,k}^{k}=1$ if the micro-batch executes on the original pipeline. We ensure that each operation is assigned only one pipeline, i.e., $\sum_{\begin{subarray}{c}k_{s}\end{subarray}}S_{i,j,k}^{k_{s}}=1$.
The MILP also takes as input profiled statistics. First, $T_{c}$ represents the computational time for each operation $c$. $T_{comm}$ represents the communication latency to send activations or gradients across workers [^1] We use a single value as activations and gradients are the same size.. Next, the MILP requires the change in memory utilization on worker $(i,k,k_{s})$ caused by executing an operation $(i,j,k,c,k_{s})$ as $\Delta M_{i,j,k,c}^{k_{s}}$.
$\Delta M_{i,j,k,c}^{k_{s}}=\begin{cases}A_{B}&\text{, if }c=F\text{ and }S_{i,j,k}^{k_{s}}=1\\ A_{B}-A_{B_{Input}}&\text{, if }c=B_{Input}\text{ and }S_{i,j,k}^{k_{s}}=1\\ -A_{B_{Weight}}&\text{, if }c=B_{Weight}\text{ and }S_{i,j,k}^{k_{s}}=1\\ 0&\text{, otherwise}\end{cases}$
Here, $A_{B}$ is the profiled size of activation stored at the end of forward pass. $A_{B_{input}}$ and $A_{B_{Weight}}$ is size of the activation stored at the end of backward-input and backward-weight pass, respectively. Upon completion of backward-input pass and backward-weight pass, we free activation $A_{B_{Input}}$ and $A_{B_{Weight}}$ respectively.
Variables. The goal of the MILP is to identify the exact ordering of operations, which can be used to generate a pipeline schedule. To do this, we introduce an ordering variable $O_{(i,j,k,c,k_{s})\rightarrow(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})}\in\{0,1\}$ between each pair of operations $(i,j,k,c,k_{s})$ and $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$. Each variable is 1 if $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled any time after $(i,j,k,c,k_{s})$ and 0 otherwise. Ordering is only required within a stage, and between computation phases of same micro-batch (e.g., between the $F$ and $B_{input}$ pass). Finally, we define the set of variables $E_{i,j,k,c}^{k_{s}}$ as the ending time of operation $(i,j,k,c,k_{s})$.
Objective. The goal of the MILP is to minimize the makespan of a single training iteration (which is repeated throughout the training process). This can be done by determining the sequence and time of operations through the training pipeline via the sets of variables $O$ and $E$. This objective can thus be formulated as:
(1) $\min_{\begin{subarray}{c}O,E\end{subarray}}\qquad\max_{\begin{subarray}{c}i,j,k,k_{s}\end{subarray}}E_{i,j,k,B_{Weight}}^{k_{s}}$
Here, $\max_{i,j,k,k_{s}}E_{i,j,k,B_{Weight}}^{k_{s}}$ represents the end time of the last operation in the iteration, representing the makespan.
Constraints. The constraints applied to the objective are the following.
Cross-Stage Dependencies.
(2) $E_{i,j,k,F}^{k_{s}}\geq S_{i,j,k}^{k_{s}}\times(\sum_{\hat{k}}(E_{i-1,j,k,F}^{\hat{k}}\times S_{i-1,j,k}^{\hat{k}})\\ +T_{comm}+T_{F})$
[…]
Equation 2 specifies the dependency of a given micro-batch’s forward pass on previous pipeline stages (e.g., stage 0 must execute before stage 1). Similarly, Equation 3 specifies the reverse dependency for the backward pass (e.g., stage 1 must execute before stage 0).
Same-Stage Dependencies.
(4) $E_{i,j,k,B_{Weight}}^{k_{s}}\geq S_{i,j,k}^{k_{s}}\times(E_{i,j,k,B_{Input}}^{k_{s}}+T_{B_{Weight}})$
Equation 4 allows the MILP to reason about Decoupled BackProp, specifying that $B_{input}$ must precede $B_{weight}$ for a given micro-batch within a worker.
No Overlapping Computations.
[…]
Equation 5 adds a dependency constraint which specifies that two different operations cannot overlap in time if they are executed on the same worker. For example, if operation $(i,j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled after $(i,j,k,c,k_{s}^{\prime})$, they both execute on worker $(i,k_{s}^{\prime})$ and thus $(i,j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ must begin after $(i,j,k,c,k_{s}^{\prime})$ ends.
Memory Constraint.
[…]
Finally Equation 6 calculates the activation memory required at any given time on worker $(i,k_{s}^{\prime})$, and constrains operations such that the total memory is below an $M_{limit}$ to avoid OOMs.
4.3. Implementation
We implemented SlipStream on top of DeepSpeed (Rasley et al., 2020) and made the following additions to DeepSpeed to support SlipStream.
Rerouting micro-batches to data parallel peers. To handle the dynamic rerouting of micro-batches following node failures, we introduce two new and complementary communication operators: ReRouteAct and ReRouteGrad. Positioned at the end of a pipeline stage, ReRouteAct normally acts as a pass-through, transmitting computed intermediate activations to the next stage within the same data parallel pipeline. Conversely, ReRouteGrad, located at the beginning of a stage, typically forwards gradients backward through the pipeline. When a subsequent stage fails, ReRouteAct redistributes the micro-batches across the remaining peers in a round-robin fashion, while ReRouteGrad adjusts the gradient distribution to ensure that both the forward and backward processes of a micro-batch are handled by the same peer. This strategy maintains operational continuity in SlipStream under fault conditions, mirroring fault-free execution semantics. These operators are integrated as pipeline instructions within the DeepSpeed execution engine.
Decoupling Back Propagation in DeepSpeed. Our implementation of decoupled backpropagation within DeepSpeed centers around intercepting the weight gradient computations traditionally performed during the backpropagation phase. These computations are temporarily held in a newly created in-memory structure known as WeightGradStore, enabling SlipStream to defer weight gradient computation. To facilitate this, two new pipeline instructions have been introduced into DeepSpeed’s execution engine: InputBackwardPass and WeightBackwardPass. These instructions are incorporated into SlipStream’s execution plan and are then processed by DeepSpeed’s execution scheduler, effectively managing the distinct phases of backpropagation.
Bypassing Optimizer Synchronizations. In DeepSpeed, each training iteration concludes with an all-reduce collective to synchronize gradients across data parallel peers, followed by a series of validation checks at each pipeline stage to ensure numerical stability, such as monitoring for gradient overflow. If any stage identifies an issue that could compromise numerical stability, the optimizer step for that iteration is omitted to avoid exacerbating the problem. Upon successful validation, the optimizer step is executed for each stage, and the training progresses to the next iteration in a synchronized manner across all stages.
Table 1. Normalized training throughput with increasing failure rates. Fault-free training represents $1.00$, hence higher is better. Bamboo ran out of memory for GPT-3 3.35B and 6.7B.
Systems GPT-3 Medium GPT-3 3.35B GPT-3 6.7B
Failure Frequency 6h 2h 30m 6h 2h 30m 6h 2h 30m
Bamboo (Thorpe et al., 2023) 0.71 0.69 0.55 OOM OOM OOM OOM OOM OOM
Oobleck (Jang et al., 2023) 0.99 0.92 0.71 0.98 0.90 0.66 0.93 0.87 0.52
SlipStream 0.99 0.92 0.81 0.98 0.95 0.85 0.97 0.91 0.66
In contrast, SlipStream employs a staggered timing approach for the optimizer steps across different pipeline stages to enhance scheduling efficiency. This method of staggering is not conducive to cross-stage synchronization for validating numerical stability prior to executing the optimizer. To accommodate staggered execution while maintaining validation integrity, numerical validations are shifted from a pre-step to a post-step process. Following the all-reduce collective, each stage immediately conducts its own local validation checks but does not wait for downstream stages to complete their validations. Each stage proceeds with its optimizer step based on its individual validation results and those of preceding stages. If a downstream stage fails its validation, a rollback is initiated across all stages before advancing to the next iteration. Notably, for many optimizers, including the widely-used AdamW (Loshchilov and Hutter, 2017), such rollbacks do not entail additional memory costs because the operations involved are arithmetically reversible. This adjustment allows SlipStream to retain the advantages of staggered operations while protecting the training process from potential instabilities.
5. Evaluation
Figure 8. Training throughput, normalized to the fault-free case, for the GPT-3 Medium and GPT-3 6.7B models over the GCP trace. In 8b and 8c, the dashed lines represent the average training throughput achieved by each system within the 6h period.
5.1. Experimental Setup
Cluster Setup. We performed real-world experiments on a cluster of 32 NVIDIA A100 80GB GPUs in Azure. Specifically, we used four Standard_NC96ads_A100_v4 instances, each with 8 GPUs, 96 vCPUs, and 880 GB of host memory. Each node has an 600 GB/s NVLink intra-node interconnect, as well as a 640 Gbps inter-node interconnect over 8 NICs. We performed scalability experiments using a simulator, which we present and validate in Section 5.3.
Baselines. We compared SlipStream to the two state-of-the-art baselines discussed in Section 2.3 – Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023) – with all of their respective optimizations enabled. We report the fault-free throughput achieved by DeepSpeed (Rasley et al., 2020) using an 1F1B pipeline schedule (Narayanan et al., 2019). All baselines train on the same Azure cluster.
Workloads. We evaluated all systems using the Megatron (Shoeybi et al., 2019) implementation of GPT-3 (Brown et al., 2020). We used three GPT-3 sizes: Medium (350M), 3.35B, and 6.7B. We used (PP, DP) degrees of (2, 16), (4, 8), and (8, 4) for the Medium, 3.35B, and 6.7B models, respectively. All models used a TP degree of 1. We trained on wikitext (Merity et al., 2016) using (batch, micro-batch) size of (8192, 8), (1024, 1), and (1024, 1) for the Medium, 3.35B, and 6.7B models, respectively. Unless otherwise noted, all real-world experiments ran for 6 hours.
5.2. Training Throughput Under Failures
How well does SlipStream handle failures compared to baselines?
For each 32-GPU training job on Azure, we instrumented a random GPU to fail at regular and increasing intervals: 6 hours, 2 hours, and 30 minutes, which is representative of real-world failure rates (Jeon et al., 2019; Weng et al., 2022), starting at the beginning of the training job ($t=0$). Note that we treat each individual GPU as a worker. Table 1 shows the average training throughput achieved by each system across models and failure rates normalized to the fault-free throughput achieved by DeepSpeed ($1.00$). Note that GPUs are not replaced after failure, hence the total number of workers monotonically declines throughout each experiment. With a failure every 30 minutes, the percentage of functional GPUs by the end of the 6 hour run is 62.5%. Hence, we expect the average throughput with a failure every 30min to be relatively modest compared to the fault-free case.
Because Bamboo relies on redundant computation, it rapidly exhausts GPU memory with its additional copy of model states and activations. Hence, Bamboo fails to train the two larger models. Even for the smallest model, redundant computation introduces significant throughput degradation. A single GPU failure (3% of total GPUs) in the 6h job causes a 29% throughput drop. Bamboo actually penalizes fault-free execution by nearly the same percentage compared to DeepSpeed. Oobleck can handle all model sizes and achieves significant throughput improvements over Bamboo. However, as the frequency of failures and model size increase, Oobleck’s performance degrades. As failures increase, Oobleck’s heterogeneous pipelines become more imbalanced, degrading the throughput of the synchronous training job. As the model size increases, the latency of reconfiguring all of the stages in the pipeline with the failed GPU gets higher.
SlipStream is able to efficiently continue training under failures, out-performing or matching the throughput of both baselines across all models and failure rates. Adaptive Pipelining, Decoupled BackProp, and Staggered Optimizer are effective in optimizing the use of compute and memory resources across peers of failed GPUs. Moreover, SlipStream recovers from a failure quickly as it migrates model parameters to only a single GPU during normalization (see Section 4.2.1). Under frequent failures (every 30m) and no repairs, SlipStream achieves 85% and 66% of the fault-free training throughput for GPT-3 3.35B and 6.7B respectively. SlipStream out-performs Oobleck and Bamboo by up to 1.29×1.29$\times$1.29 × and 1.46×1.46$\times$1.46 ×, respectively for these experiments. Note that, unlike Bamboo, both SlipStream and Oobleck introduce no penalty over DeepSpeed for fault-free execution.
What is SlipStream’s throughput advantage in an realistic training scenario?
Next, we evaluated a realistic training scenario where GPUs fail and re-join based on a real-world trace. We used a trace of node failures/repairs from GCP instances used by Bamboo (Thorpe et al., 2023). We observed similar results for Bamboo’s AWS trace, but omitted the plots due to space constraints. We replayed the events in the GCP trace over a 6h training run on our Azure cluster. At the time of the experiment, we had 24 GPUs available (instead of 32 GPUs). Figure 8 shows the number of available GPUs over time according to the trace (max 24, min 15).
Figure 8 also plots the training throughput achieved by Bamboo, Oobleck, and SlipStream replaying the trace for a GPT-3 Medium and GPT-3 6.7B training job. Solid lines represent instantaneous training throughput, while dashed lines show the average training through for each system over the 6h period. Bamboo cannot train GPT-3 6.7B due to memory pressure. SlipStream achieves a 1.64×1.64$\times$1.64 × higher performance (average throughput) compared to Bamboo on GPT-3 Medium and a 1.46×1.46$\times$1.46 × higher performance than Oobleck on GPT-3 6.7B. As discussed above, Bamboo suffers from the overheads of redundant computations. Oobleck requires large stalls due to parameter re-shuffling both when GPUs fail and when they rejoin, causing frequent and large drops in throughput. Meanwhile, during periods of stability, Oobleck suffers from imbalanced heterogeneous pipelines as discussed above. SlipStream provides the highest and most consistent training throughput across all portions of the trace.
5.3. SlipStream Scalability
Table 2. Gap between real-world and simulated throughput across various models and failure rates.
Models Fault-Free 6h 2h 30m
GPT-3 Medium -0.87% +5.98% -1.93% -1.48%
GPT-3 3.35B -0.13% -1.58% +2.12% -1.90%
GPT-3 6.7B +3.94% +2.71% -1.86% -0.85%
Since we do not have access to a cluster of thousands of GPUs, we built a simulator that can calculate the training throughput of a model and cluster configuration, given an execution plan. The simulator uses real-world profiled statistics for each pipeline operation of the respective model. To demonstrate that the simulator yields accurate results, Table 2 shows the gap between the simulated throughput and measured real-world througpught on Azure across different failure rates for the three GPT-3 models. The maximum difference is $5.98\%$. These differences arise due to minor fluctuations in the execution time of NCCL collectives, which do not cause impactful discrepancies when evaluating SlipStream.
How well does SlipStream scale to large clusters and models?
Using the simulator, we explored how well SlipStream handles failures for large models training on larger clusters. Figure 9 shows of the simulated throughput achieved by SlipStream across four different GPT-3 models with up to 145.6B parameters, normalized to the simulated fault-free throughput achieved by DeepSpeed’s 1F1B schedule. We also show the fault-scaled throughput, which is obtained by multiplying the fault-free throughput by the percent of non-failed GPUs. Note that we report the steady-state throughput with 1%, 5%, and 10% failed GPUs, not the throughput as some number of GPUs fails and rejoins. We simulated a cluster of (# GPUs, # stages per pipeline, and # pipelines) = (256, 8, 32), (512, 16, 32), (1024, 32, 32) and (1536, 64, 24) for the 18.4B, 39.1B, 76.1B, and 145.6B model, respectively, which reflect the larger cluster sizes needed to train large models.
Figure 9. Simulated throughput of SlipStream as model size increases, normalized to the simulated fault-free 1F1B throughput. The Fault-Scaled throughput is the fault-free throughput scaled by the percent of non-failed GPUs.
SlipStream is effective at achieving high throughput across failure rates and model sizes. For the case of 1% failures (higher reliability), SlipStream handles failures at or better than the fault-scaled throughput across all models and cluster sizes, often with zero degradation. The bubbles in peer workers are more than sufficient to efficiently execute the work for failed GPUs using SlipStream’s optimizations (see Figure 6). As the number of failed GPUs increases, SlipStream is still able to sustain high throughput. At $5\%$ GPU failures, SlipStream achieves comparable performance to the fault-scaled throughput. Even at a large $10\%$ failure rate (e.g., 154 failed GPUs for GPT-3 145.6B), SlipStream enables clusters to continue training while only observing between $0.5\%$ and $11.5\%$ degradation to the fault-scaled case. This suggests that SlipStream will be effective for training jobs in supercomputing-scale clusters with dynamic resource availability (Jiang et al., 2024; He et al., 2023; Zu et al., 2024).
5.4. SlipStream Performance Breakdown
Figure 10. The contribution of the three optimization techniques to SlipStream’s training throughput.
Figure 11. Memory utilization for SlipStream across pipeline stages for GPT-3 6.7B with 30m failure rates compared to DeepSpeed (failure-free). The dashed line shows the 80GB memory capacity of the A100 GPU.
How does each of SlipStream’s techniques contribute to performance?
To understand the benefit (and necessity) of each of the SlipStream techniques described in Section 3, we performed an ablation study by repeating the real-world experiment on the Azure cluster with the GPT-3 models and 30min failure frequency. We modified the MILP formulation in the Planner to progressively enable optimizations. Figure 10 shows the throughput achieved in each step of the study, normalized to the fault-free throughput.
Adaptive Pipelining allows SlipStream to continue training in the face of failures, but suffers from a large throughput degradation due to overheads caused by extra work in the peers of failed workers. By adding Decoupled BackProp, SlipStream boosts the training throughput by 63% to 118% by using bubbles in the cool-down phase of the pipeline schedule to mask this overhead (see Figure 3(b)). The addition of the Staggered Optimizer further improves throughput by 7% to 11% up to $11\%$ as it allows the use of bubbles from the warm-up phase of the pipeline schedule.
Can SlipStream efficiently exploit unused GPU memory?
SlipStream makes the key insight that GPU memory capacity is not uniformly utilized throughout the pipelined execution. Later pipeline stages inherently use less GPU memory. Decoupled BackProp utilizes this free GPU memory as a buffer to defer $B_{weight}$ computations (see Figure 5). To explore this in more detail, we analyzed the real-world training of the GPT-3 6.7B model with a 30m failure interval as in Section 5.2. Figure 11 shows the peak GPU memory utilization across the 8 stages of a pipelined execution, compared to the memory utilization of the failure-free DeepSpeed baseline.
DeepSpeed, and any system that uses a pipeline schedule similar to 1F1B, does not fully utilize GPU memory, especially in later pipeline stages (Narayanan et al., 2019). This is because GPU workers that execute earlier pipeline stages must store a larger number of intermediate results, as shown in Figure 3. Figure 11 shows how SlipStream exploits this opportunity and uniformly utilizes nearly all GPU memory to optimize pipelined execution. The MILP formulation is able to carefully model and manage the additional memory requirements within GPU memory constraints and avoid memory exhaustion.
Table 3. SlipStream Planner latency (in seconds) to find optimized schedules for up to $25\%$ of failed GPUs.
DP PP Degree
Degree 2 4 8 16 32 64
2 5.24 5.41 15.48 42.73 129.57 459.57
4 7.65 8.62 24.48 59.75 190.54 719.4
8 8.56 9.57 30.74 81.42 294.21 1038.2
16 11.84 13.8 51.53 116.8 452.98 1721.45
32 11.71 14.87 67.24 278.64 942.42 3153.07
What is the overhead of SlipStream’s Planner?
To understand how fast the Planner can generate optimized schedules for large clusters, Table 3 shows the time required to generate all necessary schedules for up to $25\%$ of node failures across various parallelism strategies. Recall that the user can configure a smaller or larger fault-tolerance threshold. The reported latency includes both Planner phases. We report the results using Gurobi MILP solver (Gurobi, 2024) running on an 96-core CPU. We can generate all solutions to support up to $512$ failures in the 2048-GPU case ($DP=32$, $PP=64$) within 3153 seconds (52.5 minutes). Given that the training time for large models on thousands of GPUs is measured in weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020), the Planner’s latency is insigificant.
6. Related Work
Parallel Training. Data parallelism (Krizhevsky et al., 2017; Chen et al., 2016; Dean et al., 2012) is a widely utilized mode of parallelism that distributes the dataset across different partitions for processing. In this mode, the learned weights are synchronized via either an all-reduce approach (Chen et al., 2016) or by using parameter servers (Li et al., 2014; Chilimbi et al., 2014; Jiang et al., 2020). Alternatively, model parallelism (Krizhevsky, 2014; Shoeybi et al., 2019; Shazeer et al., 2018) involves distributing the components of a deep neural network (DNN) model across multiple GPU devices, allowing each device to handle a specific subset of the model’s parameters for all input data. Recently, pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023) has emerged as a technique for training large models by dividing the model’s layers among different workers and utilizing microbatches to efficiently use the available computational resources. Prominent deep learning training frameworks like PyTorch (PyTorch, 2024), DeepSpeed (Rasley et al., 2020), and Megatron (Narayanan et al., 2021) have adopted hybrid-parallelism, an approach that integrates data parallelism, model parallelism, and pipeline parallelism. This integration facilitates training on a massive scale while enhancing computational and memory efficiency. Additionally, DeepSpeed introduces ZeRO-style (Rajbhandari et al., 2020, 2021; Ren et al., 2021; Zhao et al., 2023; Korthikanti et al., 2023) data parallelism, which strategically partitions model states across GPUs, coordinating via communication collectives to synchronize parameters as needed.
Optimizing Hybrid-Parallel Training. Improvements in hybrid parallelism are extensively studied in the context of ML training (Narayanan et al., 2021; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022; Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023). Alpa (Zheng et al., 2022) automatically optimizes inter- and intra-operator parallelism using a hierarchical ILP formulation. Tofu (Wang et al., 2019a) uses DP to optimally partition tensor operations in a single node. FlexFlow (Jia et al., 2018) uses a randomized search algorithm to quickly find parallelism strategies. TensorOpt (Cai et al., 2022) introduces a DP that can optimize parallelism for multiple resource dimensions (e.g., memory and compute). Piper (Tarnawski et al., 2021) proposes a two-level DP algorithm to find optimal hybrid parallelism strategies. SlipStream uses a DP algorithm and an MILP which exploit key functional redundancies in hybrid parallelism to explicitly improve training throughput under failures.
Elastic Training. Elastic training systems can dynamically adjust the resources allocated to a training job. Both Horovod (Sergeev and Del Balso, 2018) and Torch Distributed (Li et al., 2020) provide mechanisms to modify the number of workers, but require either a restart from a checkpoint or an expensive re-shuffling of model parameters. CoDDL (Hwang et al., 2021), Optimus (Peng et al., 2018), OASiS (Bao et al., 2018), and Themis (Mahajan et al., 2020) are ML cluster schedulers that can dynamically allocate resources across multiple DNN training jobs. Or et al. (Or et al., 2020) auto-scale the number of workers for a training job. Varuna (Athlur et al., 2022) provides elastic training by leveraging spot instances, but requires restarts from checkpoints to handle unexpected pre-emptions. While some elastic training systems can affect model consistency by changing important hyperparameters such as batch size and learning rate, SlipStream ensures that operations are mathematically consistent regardless of the number of failures.
Fault-tolerant Training. DNN training systems widely leverage checkpoints as the basic mechanism for fault recovery (Rasley et al., 2020; Zu et al., 2024; Le Scao et al., 2022; Jiang et al., 2024; He et al., 2023; Beaumont, 2022; Jeon et al., 2019; Wang et al., 2019b). While naive checkpointing can lead to training stalls, recent work aims to reduce these overheads. CheckFreq (Mohan et al., 2021) dynamically adapts the checkpointing frequency to reduce overheads. Check-N-Run (Eisenman et al., 2022) quantizes in-frequently updated embedding table checkpoints for recommendation models. Gemini (Wang et al., 2023) introduces checkpoint traffic scheduling across the storage hierarchy (local CPU memory, remote CPU memory, and remote persistent storage). Megascale (Jiang et al., 2024) alleviates the storage bottleneck during recovery by sharing data between corresponding GPU workers across data parallel groups.
Two recent projects have proposed to utilize pipeline parallelism for efficient training in the presence of failures without spares. Bamboo (Thorpe et al., 2023) introduces redundant computation (RC) to provide resilience in the presence of frequent preemptions of training with spot instances. Oobleck (Jang et al., 2023) provides resiliency through creation of heterogenous pipelines. SlipStream matches or outperforms them for a wide range of model sizes and failure frequencies.
7. Conclusion
SlipStream enables efficient DNN training in the presence of failures without the use of spare resources. SlipStream exploits the functional redundancies inherent to hybrid parallel training systems used for large DNNs. It introduces novel techniques to minimize the training throughput degradation of sustaining failures by exploiting unused resources (e.g., pipeline bubbles). We evaluated an end-to-end SlipStream prototype and showed it can tolerate a high numbers of concurrent failures for systems and models of various sizes. SlipStream improves training throughput under failures over Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 × respectively.

### v2

@@
-SlipStream: Adapting Pipelines for Distributed Training of Large DNNs Amid Failures
+ReCycle: Resilient Training of Large DNNs using Pipeline Adaptation
 Abstract.
-Training large Deep Neural Network (DNN) models requires thousands of GPUs for days or weeks at a time.
-At these scales, failures are frequent and can have a big impact on training throughput.
-Restoring performance using spare GPU servers becomes increasingly expensive as models grow.
-SlipStream is a system for efficient DNN training in the presence of failures, without using spare servers.
-It exploits the functional redundancy inherent in distributed training systems – servers hold the same model parameters across data-parallel groups – as well as the bubbles in the pipeline schedule within each data-parallel group.
-SlipStream dynamically re-routes the work of a failed server to its data-parallel peers, ensuring continuous training despite multiple failures.
-However, re-routing work leads to imbalances across pipeline stages that degrades training throughput.
-SlipStream introduces two optimizations that allow re-routed work to execute within bubbles of the original pipeline schedule.
-First, it decouples the backward pass computation into two phases.
-Second, it staggers the execution of the optimizer step across pipeline stages.
-Combined, these optimizations enable schedules that minimize or even eliminate training throughput degradation during failures.
-We describe a prototype for SlipStream and show that it achieves high training throughput under multiple failures, outperforming recent proposals for fault-tolerant training such as Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 ×, respectively.
-copyright: none
+Training large Deep Neural Network (DNN) models requires thousands of GPUs over the course of several days or weeks.
+At this scale, failures are frequent and can have a big impact on training throughput.
+Utilizing spare GPU servers to mitigate performance loss becomes increasingly costly as model sizes grow.
+ReCycle is a system designed for efficient DNN training in the presence of failures, without relying on spare servers.
+It exploits the inherent functional redundancy in distributed training systems – where servers across data-parallel groups store the same model parameters – and pipeline schedule bubbles within each data-parallel group.
+When servers fails, ReCycle dynamically re-routes micro-batches to data-parallel peers, allowing for uninterrupted training despite multiple failures.
+However, this re-routing can create imbalances across pipeline stages, leading to reduced training throughput.
+To address this, ReCycle introduces two key optimizations that ensure re-routed micro-batches are processed within the original pipeline schedule’s bubbles.
+First, it decouples the backward pass into two phases: one for computing gradients for the input and another for calculating gradients for the parameters.
+Second, it avoids synchronization across pipeline stages by staggering the optimizer step.
+Together, these optimizations enable adaptive pipeline schedules that minimize or even eliminate training throughput degradation during failures.
+We describe a prototype for ReCycle and show that it achieves high training throughput under multiple failures, outperforming recent proposals for fault-tolerant training such as Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 ×, respectively.
+Fault-tolerant Training, Distributed Training, Hybrid Parallelism, Pipeline Adaptation
+ccs: Computer systems organization Dependable and fault-tolerant systems and networksccs: Computer systems organization Distributed architecturesccs: Computer systems organization Neural networkscopyright: acmlicensedjournalyear: 2024conference: ACM SIGOPS 30th Symposium on Operating Systems Principles; November 4–6, 2024; Austin, TX, USAbooktitle: ACM SIGOPS 30th Symposium on Operating Systems Principles (SOSP ’24), November 4–6, 2024, Austin, TX, USAdoi: 10.1145/3694715.3695960isbn: 979-8-4007-1251-7/24/11
 1. Introduction
@@
 Bamboo’s drawback is the drop in fault-free training throughput due to the redundant computations and added memory pressure on each GPU node.
-Oobleck (Jang et al., 2023) avoids overheads in fault-free execution by pre-computing a number of templates for pipeline parallelism that use a different number of nodes.
+Oobleck (Jang et al., 2023) avoids overheads in fault-free execution by precomputing a number of templates for pipeline parallelism that use a different number of nodes.
 As nodes fail, Oobleck replaces the original pipeline template with one that uses fewer nodes.
@@
 Both Bamboo and Oobleck require significant re-configuration as nodes fail or rejoin, which can be a challenge as failure rates increase.
-In this paper, we introduce SlipStream, a novel scheme for resilient training with hybrid-parallel systems.
-SlipStream enables efficient execution in the presence of failures, without the need for spares and without any impact on model accuracy.
+In this paper, we introduce ReCycle, a novel scheme for resilient training with hybrid-parallel systems.
+ReCycle enables efficient execution in the presence of failures, without the need for spares and without any impact on model accuracy.
 Similar to Bamboo and Oobleck, we exploit pipeline parallelism.
-But unlike these previous proposals that handle failures within a single pipeline, SlipStream utilizes the inherent functional redundancy across pipelines in hybrid-parallel training systems.
-SlipStream re-routes the micro-batches of failed nodes to peer nodes that process the same pipeline stage for the model in the other data-parallel groups.
+But unlike these previous proposals that handle failures within a single pipeline, ReCycle utilizes the inherent functional redundancy across pipelines in hybrid-parallel training systems.
+ReCycle re-routes the micro-batches of failed nodes to peer nodes that process the same pipeline stage for the model in the other data-parallel groups.
 Peer nodes store the same model parameters and can handle the extra work without the need for parameter re-shuffling.
-If done naively, SlipStream’s adaptive pipelining increases training time and memory usage due to the additional work for the peers of failed nodes.
-SlipStream uses a series of optimizations that exploit unique characteristics of hybrid-parallel training to eliminate all or most of the inefficiencies.
-SlipStream uses the bubbles in the pipeline schedule of peers to execute the additional work at low or no overhead.
+If done naively, ReCycle’s adaptive pipelining increases training time and memory usage due to the additional work for the peers of failed nodes.
+ReCycle uses a series of optimizations that exploit unique characteristics of hybrid-parallel training to eliminate all or most of the inefficiencies.
+ReCycle uses the bubbles in the pipeline schedule of peers to execute the additional work at low or no overhead.
 These bubbles typically appear during the start-up and cool-down phase of the pipeline schedule, but not during the steady-state when most additional work must occur.
-SlipStream overcomes this challenge by decoupling back propagation for micro-batches into two distinct gradient computations, one relative to the input and one relative to the parameters.
+ReCycle overcomes this challenge by decoupling back propagation for micro-batches into two distinct gradient computations, one relative to the input and one relative to the parameters.
 The two steps are independently scheduled in order to better leverage bubbles in the the cool-down portion of the pipeline schedule.
 The selective application of decoupled back propagation exploits the imbalance in memory usage across stages in hybrid-parallel training while avoiding an increase in peak memory pressure.
-Finally, SlipStream staggers the execution of the optimizer tasks across pipeline stages in order to leverage bubbles from the start-up phase of the pipeline schedule for the next training iteration.
-We implemented SlipStream on top of the DeepSpeed training framework (Rasley et al., 2020).
-The key SlipStream element is the Planner that uses dynamic programming (DP) and mixed integer linear programming (MILP) to find the best pipeline schedule in the presence of multiple failures.
-The Planner runs offline and calculates efficient schedules for up to a certain number of failures.
-These schedules are employed as needed during training.
-We evaluated SlipStream using DNNs with billions of parameters like GPT-3 (Brown et al., 2020), using both real-world experiments and simulations of large training systems.
-We show that SlipStream supports high throughput training even at high failure counts, e.g.
+Finally, ReCycle staggers the execution of the optimizer tasks across pipeline stages in order to leverage bubbles from the start-up phase of the pipeline schedule for the next training iteration.
+We implemented ReCycle on top of the DeepSpeed training framework (Rasley et al., 2020).
+The key component of ReCycle is the Planner, which utilizes dynamic programming and mixed-integer linear programming (MILP) to determine adaptive pipeline schedule in presence of multiple failures.
+Operating offline, the Planner precomputes efficient schedules for a predefined number of failures, which are then applied as necessary during training.
+We evaluated ReCycle using DNNs with billions of parameters like GPT-3 (Brown et al., 2020), using both real-world experiments and simulations of large training systems.
+We show that ReCycle supports high throughput training even at high failure counts, e.g.
 $\geq$10of the overall system.
-SlipStream outperforms Oobleck by 1.46×1.46$\times$1.46 × in training scenarios with realistic GPU failures.
-SlipStream’s advantage stems from its efficient scheduling of re-routed work and its ability to avoid major re-shuffling of model parameters upon failures.
-SlipStream outperforms Bamboo by 1.64×1.64$\times$1.64 ×.
+ReCycle outperforms Oobleck by 1.46×1.46$\times$1.46 × in training scenarios with realistic GPU failures.
+ReCycle’s advantage stems from its efficient scheduling of re-routed work and its ability to avoid major re-shuffling of model parameters upon failures.
+ReCycle outperforms Bamboo by 1.64×1.64$\times$1.64 ×.
 It is also able to efficiently train much larger models than Bamboo, which requires large memory overheads for its fault tolerance mechanism.
@@
 2.1. Distributed DNN Training
-Figure 1.
-Illustration of hybrid parallelism.
-Pipeline stages are denoted with different colors.
-Within each pipeline stage, operators are partitioned through tensor parallelism.
-The global batch is split into micro-batches across pipelines.
 State-of-the-art Deep Neural Network (DNN) models consist of tens to hundreds of billions of parameters and are trained on datasets with many trillions of tokens (Brown et al., 2020; Achiam, 2023; Touvron et al., 2023; Le Scao et al., 2022; Zhang et al., 2022; Fedus et al., 2021; Lepikhin et al., 2020).
@@
 However, it achieves lower compute utilization as the number of stages increases due to pipeline dependencies and bubbles (idle slots) in the pipeline schedule.
-Large-scale training systems balance these tradeoffs by utilizing all three forms of parallelism as shown in Figure 1 (Narayanan et al., 2021; Li and Hoefler, 2021; Liu et al., 2023; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022).
+Large-scale training systems balance these trade-offs by utilizing all three forms of parallelism as shown in Figure 1 (Narayanan et al., 2021; Li and Hoefler, 2021; Liu et al., 2023; Korthikanti et al., 2023; Rasley et al., 2020; Li et al., 2020; Zhao et al., 2023; Zheng et al., 2022).
 This is known as hybrid-parallelism.
@@
 Hybrid parallelism enables large model training with graceful scaling and reasonable training times (weeks to few months) using optimized clusters with 1000s of densely-connected GPUs.
+Figure 1.
+Illustration of hybrid parallelism.
+Pipeline stages are denoted with different colors.
+Within each pipeline stage, operators are partitioned through tensor parallelism.
+The global batch is split into micro-batches across pipelines.
 2.2. Fault Tolerance in Distributed DNN Training
@@
 At this scale, faults are regular occurrences that require systematic optimizations to ensure resilient and efficient training (He et al., 2023; Jiang et al., 2024; Jeon et al., 2019; Weng et al., 2022).
-Enhancing fault tolerance in distributed training involves three key issues: fast error detection, efficient checkpointing, and robust execution under faults.
-Fast error detection is crucial for minimizing downtime and preventing the spread of errors.
-Some errors can be quickly detected through encoding and redundancy mechanisms in the hardware (e.g., error correction codes) or software timeouts.
-Other errors are detected if they eventually lead to irregular spikes in the loss function optimized during training (Beaumont, 2022).
-The latter category includes silent data corruptions (Bacon, 2022; Dattatraya Dixit et al., 2022).
-Checkpointing, which saves the state of the model at regular intervals, allows for recovery from the recent stable state without restarting the entire training from scratch.
-Since naive checkpointing to remote storage leads to long pauses in the training process, there has been considerable research in improving its efficiency (Mohan et al., 2021; Eisenman et al., 2022; Wang et al., 2023; Jiang et al., 2024) as discussed in Section 6.
-2.3. Efficient DNN Training in Presence of Failures
-Our work focuses on robust training in the presence of faults.
-A common approach is to have a reserve of spares ready to replace servers with failed GPU servers.
-To reduce recovery time, spare servers can also be hot standbys with their memory loaded with model parameters at the end of each iteration.
-The key advantage of spare servers is that, upon recovery, training efficiency returns to its pre-failure state.
-However, this approach becomes increasingly costly as systems scale up, fault frequency increases, and more spares are needed.
-Spare servers must connect to the network of the training cluster with a similar bandwidth profile as the servers they will be called to replace, which incurs additional networking costs and typically requires that spares are allocated in coarse-grained groups (Jouppi et al., 2023).
-Hot standby servers should be provisioned according to the full capacity of a data parallel group, determined by multiplying pipeline parallelism ($PP$) by tensor parallelism ($TP$).
-For instance, training a 530B GPT model requires 280 spare GPUs on standby, which results in an 11% increase in training costs (Narayanan et al., 2021).
-Spare servers can potentially run small training jobs while on (non hot) standby to amortize cost.
-However, the job eviction latency needed to free these resources upon a failure can cause long stalls for large training jobs occupying thousands of GPUs.
-The alternative approach is to continue training with the largest possible subset of functional resources.
-Ideally, when $x\%$ of GPU servers fail, training can continue at no less than $(100-x)\%$ of the fault-free throughput.
-The simplest method, sometimes referred to as elastic batching, is to drop a whole data parallel (DP) group when a node fails (Wang et al., 2018; Lin et al., 2019).
-The remaining data parallel groups can proceed with the same parallelization scheme.
-While straightforward, this method has substantial drawbacks.
-A single node failure results in $PP\times TP$ nodes going offline, decreasing the total throughput by $1/DP$ (e.g., 17% when training a 1000B parameter GPT model (Narayanan et al., 2021)).
-Moreover, this requires reducing the training batch size or recompiling the program to optimize each data parallel group for a larger portion of the batch.
-This can disrupt the balance and efficiency of the entire training job.
-Figure 2. Adaptive pipelining when $W_{1\textunderscore 2}$ fails. The output from worker $W_{1\textunderscore 1}$ is redirected to workers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$.
+Enhancing fault tolerance in distributed training requires addressing three critical areas: fast error detection, checkpointing, and efficient execution in presence of faults.
+2.2.1. Error Detection
+Fast Error Detection is vital for minimizing downtime and preventing the propagation of errors.
+Effective error detection mechanisms include both hardware and software solutions.
+Hardware mechanisms, such as error correction codes (ECC), provide immediate detection and correction of data corruption.
+Software solutions can involve timeouts and heartbeat signals to identify system failures quickly.
+Additionally, some errors, particularly those that are silent or not immediately apparent, are detected through irregularities in the loss function.
+These anomalies, often indicative of silent data corruptions, can significantly impact training outcomes if not addressed promptly (Beaumont, 2022; Bacon, 2022; Dattatraya Dixit et al., 2022).
+2.2.2. Checkpointing
+Checkpointing plays a crucial role in fault tolerance by allowing the system to recover from a recent stable state rather than restarting the entire training process.
+It involves saving the model’s state at regular intervals.
+However, naive checkpointing methods that write data to remote storage can introduce significant pauses in the training process due to latency and bandwidth constraints.
+To address this, extensive research has focused on optimizing checkpointing techniques to reduce overhead and improve efficiency (Mohan et al., 2021; Eisenman et al., 2022; Wang et al., 2023; Jiang et al., 2024), as detailed in Section 7.
+2.2.3. Efficient DNN Training in the Presence of Failures
+Our work focuses on ensuring that the system remains operational even when parts of it fail.
+A common method is to use warm spares, preloaded with model parameters to reduce recovery time.
+While this ensures that training efficiency returns to pre-failure levels, it becomes costly as systems scale and fault frequency rises.
+Spare servers need similar network bandwidth as the ones they replace, adding networking costs and requiring coarse-grained allocation (Jouppi et al., 2023).
+For example, training a 530B GPT model requires 280 spare GPUs, provisioned according to the full capacity of a data parallel group, increasing costs by 11% (Narayanan et al., 2021).
+Use of cold spares to run small jobs can offset some of these costs, however eviction delays upon failure can cause significant stalls for large jobs.
+Figure 2.
+Adaptive pipelining when $W_{1\textunderscore 2}$ fails.
+The micro-batches from worker $W_{1\textunderscore 1}$, originally intended for $W_{1\textunderscore 2}$, are dynamically re-routed to workers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$, ensuring that the training process continues without interruption.
 (a) Fault-Free 1F1B Schedule
@@
 Figure 3.
-Hybrid-parallel training with 3 data-parallel pipelines, 4 pipeline stages, and 6 micro-batches per iteration.
+Hybrid-parallel training across 12 workers with 3 data-parallel pipelines, each with 4 pipeline stages, and 6 micro batches.
 3(a) shows a 1F1B fault-free schedule.
-3(b) shows how Adaptive Pipelining re-routes micro-batches failed workers to functional peers.
-Two recent projects have instead utilized pipeline parallelism for efficient training in the presence of failures without spares.
-Bamboo (Thorpe et al., 2023) draws inspiration from RAID (Patterson et al., 1988) and designs a pipeline for redundant computation (RC).
-Each pipeline stage is redundantly computed on two nodes, even in the fault-free case.
-When a node fails, its backup node steps up and computes the forward and backward passes of the failed node.
-Bamboo uses RC even in the absence of faults.
-While RC tries to hide some overhead in pipeline bubbles, its extra computational overhead still significantly drops training throughput (see Section 5).
-RC also replicates the state of each pipeline stage in two nodes, increasing GPU memory pressure and limiting its scalability.
-Finally, Bamboo needs to restart with a full reconfiguration from a checkpoint for as few as two failures, e.g., when two adjacent nodes fail.
-Oobleck (Jang et al., 2023) adapts pipeline parallelism for resilient execution but has no overhead in the fault-free case.
-It relies on pre-computed templates for pipeline parallelism.
-All templates are logically equivalent yet physically heterogeneous.
-Each template has a different number of stages, a different micro-batch configuration, and occupies a different number of nodes.
-A pipeline with failed node(s) is replaced with a new pipeline from a template that requires fewer nodes.
-However, templates with fewer nodes have lower training throughput.
-To avoid having the slowest pipeline slowdown the overall synchronous training job, Oobleck attempts to distribute the global mini-batch proportionally to compute power.
-This is not always possible and the straggling pipelines will inflate training time.
-Moreover, each pipeline re-configuration is in the critical path of training and can become an overhead if nodes fail frequently.
-Re-configuration is also needed when nodes re-join.
-Finally, Oobleck treats each heterogeneous pipeline as a black-box.
-It fails to observe that there are bubbles across all pipelines that can help hide some or all of the overheads due to node failures.
-3. SlipStream Techniques
-SlipStream aims to support efficient distributed training in the presence of faults.
+3(b) shows how Adaptive Pipelining re-routes micro-batches from failed worker $W_{1\textunderscore 2}$ to its functional peers $W_{0\textunderscore 2}$ and $W_{2\textunderscore 2}$.
+An alternative approach is to continue training with the largest subset of available resources.
+Ideally, if $x\%$ of GPU servers fail, training proceeds at no less than $(100-x)\%$ of fault-free throughput.
+A simple method, known as elastic batching, drops an entire data parallel group when a node fails, allowing the remaining groups to continue with the same parallelization (Wang et al., 2018; Lin et al., 2019).
+However, this approach has significant drawbacks, as a single node failure can take $PP\times TP$ nodes offline, reducing throughput by $1/DP$ (e.g., 17% for a 1T parameter GPT model (Narayanan et al., 2021)).
+Moreover, this approach requires reducing the batch size or recompiling the program to optimize each data parallel group for a larger batch portion, disrupting the balance and efficiency of the training job.
+Two recent projects use pipeline parallelism for fault-tolerant training without spares.
+Bamboo (Thorpe et al., 2023), drawing inspiration from RAID (Patterson et al., 1988), employs redundant computation (RC) where each pipeline stage is replicated on two nodes, even in fault-free cases.
+When a node fails, its backup handles its and forward and backward passes of failed node.
+Though RC hides some overhead in pipeline bubbles, it still significantly reduces throughput (see Section 6) and increases GPU memory pressure, limiting scalability.
+Moreover, Bamboo must restart with a full reconfiguration from a checkpoint for as few as two adjacent node failures.
+Oobleck (Jang et al., 2023) uses pipeline parallelism to ensure resilient execution with no overhead when no faults occur.
+It utilizes precomputed templates for pipeline parallelism, each differing in stages, micro-batch configurations, and node counts.
+If a pipeline encounters a failed node, it switches to a template with fewer nodes, which may reduce training throughput.
+To prevent a slow pipeline from affecting the overall job, Oobleck distributes the global mini-batch based on compute power, leading to increased training time.
+Pipeline re-configuration when nodes fail or re-join can also become an overhead if failures are frequent.
+Additionally, Oobleck treats each pipeline as a black-box and does not leverage bubbles to mitigate overheads from node failures.
+3. ReCycle Techniques
+ReCycle aims to support efficient distributed training in the presence of faults.
 It requires no spare servers and has no impact on model accuracy compared to fault-free training.
-SlipStream can tolerate multiple hardware failures and maintains training throughput proportional to the number of functional servers available.
-SlipStream is optimized for fast recovery as failures do not require significant re-shuffling of model parameters between functional nodes.
-This section reviews the key SlipStream techniques: Adaptive Pipelining, Decoupled BackProp, and a Staggered Optimizer. Section 4 presents the SlipStream system design.
+ReCycle can tolerate multiple hardware failures and maintains training throughput proportional to the number of functional servers available.
+ReCycle is optimized for fast recovery as failures do not require significant re-shuffling of model parameters between functional nodes.
+This section reviews the key ReCycle techniques: Adaptive Pipelining, Decoupled BackProp, and a Staggered Optimizer. Section 4 presents the ReCycle system design.
 3.1. Adaptive Pipelining: Working Around Failures
-SlipStream exploits two key properties in hybrid-parallel training systems.
+ReCycle exploits two key properties in hybrid-parallel training systems.
 First, there is functional redundancy across data-parallel pipelines.
@@
 Adaptive Pipelining exploits functional redundancy and bubbles by dynamically re-routing micro-batches from a failed worker to its functioning peers in other data-parallel pipelines.
-We aim to use the bubbles in peers’ schedules to process the micro-batches for the failed worker at low performance penalty.
+We aim to use the bubbles in peers’ schedules to process the micro-batches for the failed worker with a low performance penalty.
 We evenly distribute micro-batches across all functional peers.
@@
 The overall mathematical computation remains unchanged from the fault-free 1F1B schedule, ensuring that Adaptive Pipelining does not impact model convergence.
-Figure 4. Forward and Backward pass for an operator.
-The scheduling challenge: In theory, the execution schedule of the two peer workers has enough bubbles to do the computations for failed node $W_{1\_2}$ as there are 18 idle slots in total in Figure 3(a).
-However, these bubbles are concentrated in the warm-up and cool-down phase of the 1F1B schedule.
+Are sufficient bubbles available?
+Adaptive Pipelining exploits the $3\times(PP-1)\times DP$ existing but previously unused bubbles in each pipeline to recover from failures.
+For example, the 405 billion parameter LLaMA-3 is trained with synchronous 1F1B over $8192$ GPUs using hybrid parallelism with tensor, pipeline, and data parallelism degree $TP=8$, $PP=16$, and $DP=64$, respectively (AI, 2024).
+ReCycle can leverage $(3\times(16-1)\times 64)=2880$ bubbles per iteration to accommodate $\frac{2880}{3}=960$ rerouted micro-batches per iteration.
+For a training job with global batch size of 2048, this is sufficient to handle 30 simultaneous failures.
+The scheduling challenge: While sufficient bubbles exist, these bubbles are concentrated in the warm-up and cool-down phase of the 1F1B schedule.
+For example, there are 18 idle slots across data-parallel peers of failed node $W_{1\_2}$ in Figure 3(a) to accommodate its micro-batches.
 Adaptive Pipelining needs to reroute micro-batches to peer workers $W_{0\_2}$ and $W_{2\_2}$ mostly in the middle of their steady-state schedule, which is optimized to be bubble-free for efficiency (Narayanan et al., 2019).
 As shown in Figure 3(b), this has a significant performance impact.
-The additional micro-batches to peer workers and the dependencies between forward and backward pass computations lead to 36 time steps per iteration.
+The additional micro-batches to peer workers and the dependencies between forward and backward pass computations lead to 9 additional time steps per iteration, for a total of 36.
 In other words, the failure of 8.3% of workers (1 out of 12) leads to a 33% slowdown in iteration time.
 3.2. Decoupled BackProp: Filling Unused Bubbles
+Figure 4. Forward and Backward pass for an operator.
 Figure 5. Optimized schedule with Decoupled BackProp when worker $W_{1\textunderscore 2}$ fails.
+(a) Fault-free 1F1B Schedule
+(b) Schedule for Adaptive Pipelining with Decoupled BackProp when worker $W_{1\textunderscore 2}$ fails.
+(c) Schedule with the addition of the Staggered Optimizer.
+Figure 6. The Staggered Optimizer allows ReCycle to optimize pipeline schedule across training iterations.
 Decoupled BackProp addresses the scheduling problem of Adaptive Pipelining.
@@
 Thus, Decoupled BackProp splits the backward pass into two distinct tasks.
-(a) Fault-free 1F1B Schedule
-(b) Schedule for Adaptive Pipelining with Decoupled BackProp when worker $W_{1\textunderscore 2}$ fails.
-(c) Schedule with the addition of the Staggered Optimizer.
-Figure 6. The Staggered Optimizer allows SlipStream to optimize pipeline schedule across training iterations.
 Figure 5 explains how Decoupled BackProp reduces the overheads introduced by Adaptive Pipelining.
@@
 Hence, we can take advantage of idle slots in the cool-down phase of the 1F1B schedule for $B_{Weight}$ computations, freeing slots in the steady stage for the re-routed work of the failed worker.
-This enables SlipStream to absorb the additional computational demands imposed by failures with minimal overheads.
+This enables ReCycle to absorb the additional computational demands imposed by failures with minimal overheads.
 The memory challenge.
@@
 For instance, as shown in Figure 5, decoupling the backward pass for micro batch 1 necessitates retaining its intermediate data until time step 25.
-To avoid memory exhaustion, SlipStream applies Decoupled BackProp selectively only when it can mitigate the overheads of Adaptive Pipelining.
+To avoid memory exhaustion, ReCycle applies Decoupled BackProp selectively only when it can mitigate the overheads of Adaptive Pipelining.
 We also capitalize on the observation that memory imbalance exists among the pipeline stages (Kim et al., 2023).
 To ensure that all pipeline stages are fully utilized, earlier stages need to allocate additional memory to process more forward micro-batches than later stages (Narayanan et al., 2019).
-By exploiting this available surplus memory, SlipStream can effectively offset some of the memory demands of Decoupled BackProp without incurring additional costs.
+By exploiting this available surplus memory, ReCycle can effectively offset some of the memory demands of Decoupled BackProp without incurring additional costs.
 3.3. Staggered Optimizer: Accessing More Bubbles
@@
 The Staggered Optimizer thus shifts the timing of the optimizer step across pipeline stages to better utilize the bubbles that occur during the warm-up phase of the subsequent training iteration.
-Effectively, this staggering provides SlipStream more bubbles in the cool-down phase to hide the overhead of compensating for failed workers.
+Effectively, this staggering provides ReCycle more bubbles in the cool-down phase to hide the overhead of compensating for failed workers.
 Figure 6(c) shows how combining Staggered Optimizer with Adaptive Pipelining and Decoupled BackProp results in zero overhead over the fault-free 1F1B schedule in the running example.
@@
 Worker $W_{0\_0}$ can start iteration 2 at exactly the same time step as the fault-free 1F1B schedule, even with failures.
-Put together, these three mechanisms allow SlipStream to flexibly optimize the use of computational resources across pipelines, ensuring that each stage operates near its capacity and maintains consistent utilization throughout training.
-3.4. Supporting Multiple Failures
-SlipStream’s techniques support multiple failures.
-SlipStream guarantees continued training up to $DP-1$ simultaneous worker failures, where $DP$ is the number of data-parallel pipelines, as one functional worker exists for each pipeline stage.
-SlipStream can probabilistically sustain more failures, as long as there is at least one functional worker per stage across all data parallel pipelines.
-Figure 7(b) shows how SlipStream can recover from more than $DP-1$ failures.
+Put together, these three mechanisms allow ReCycle to flexibly optimize the use of computational resources across pipelines, ensuring that each stage operates near its capacity and maintains consistent utilization throughout training.
+3.4. Supporting Multiple Failures and Re-Joins
+(a) The loss of an entire DP group requires checkpoint restoration.
+(b) ReCycle can continue training if a functional worker exists per stage.
+Figure 7. Example of ReCycle’s fault tolerance guarantees.
+ReCycle supports multiple failures and rejoins.
+ReCycle guarantees continued training up to $DP-1$ simultaneous worker failures, as in the worst case $DP$ failures can disable all the data-parallel peers of a pipeline stage.
+However, ReCycle can probabilistically sustain more failures, as long as there is at least one functional worker per stage across all data parallel pipelines.
+Figure 7(b) shows how ReCycle can recover from more than $DP-1$ failures.
 Despite 8 simultaneous failures – 2/3 of the GPUs!
-– SlipStream continues training.
-(a) The loss of an entire DP group requires checkpoint restoration.
-(b) SlipStream can continue training if a functional worker exists per stage.
-Figure 7. Example of SlipStream’s fault tolerance guarantees.
-As workers are repaired, SlipStream can re-insert them back into the adapted pipeline schedule, reducing the probability of high counts of concurrent failures.
-In the rare event that simultaneous failures affect a single data parallel group as shown in Figure 7(a), SlipStream resorts to current resilience solutions: calculate an efficient hybrid-parallel scheme for the remaining nodes, configure it across all workers by restoring from a recent checkpoint, and resume training at the speed the new parallelism scheme allows.
+– ReCycle ensures continual training.
+As workers are repaired, ReCycle can re-insert them back into the adapted pipeline schedule, reducing the probability of high counts of concurrent failures.
+Unlike unexpected failures, planned worker additions occur at iteration boundaries to overlap any data copying overhead with the previous iteration.
+Once a worker rejoins, ReCycle ceases rerouting micro-batches and sends all micro-batches from the previous stage to the recovered worker.
+In the worst-case scenario, where simultaneous failures impact a single data-parallel group (as shown in Figure 7(a)), ReCycle falls back on existing resilience strategies: it calculates an efficient hybrid-parallel scheme for the remaining nodes, configures it across all workers by restoring from a recent checkpoint, and resumes training at the speed supported by the new parallelism.
 Multi-GPU Servers and Tensor Parallelism.
-Training systems commonly employ servers with multiple GPUs.
-A favoured configuration involves HGX/DGX servers with 8 GPUs connected in an all-to-all manner using NVLink and NVSwitches (DGX, 2024).
-Similar to Oobleck and Bamboo, we consider a whole server as the unit of failure.
+Training systems commonly employ servers equipped with multiple GPUs.
+A favoured configuration features HGX/DGX servers with 8 GPUs connected in an all-to-all manner using NVLink and NVSwitches (DGX, 2024).
+Similar to Oobleck and Bamboo, we treat an entire server as the unit of failure.
 This approach is pragmatic because servicing even one faulty GPU board necessitates taking the entire server offline.
-Moreover, software faults, CPU issues, connectivity problems, and power or cooling issues can also incapacitate a whole server.
-We also need to account for the level of tensor parallelism.
-Most training systems apply tensor parallelism within a multi-GPU server to leverage the high-bandwidth NVLink connections.
-In cases where tensor parallelism spans multiple servers, SlipStream, along with Oobleck and Bamboo, identifies the collective group of tensor-parallel servers as a unit of failure.
-4. SlipStream Design
+Additionally, issues such as software malfunctions, CPU issues, connectivity disruptions, and power or cooling issues can incapacitate an entire server.
+Most training systems implement tensor parallelism within a multi-GPU server to leverage the high-bandwidth NVLink connections.
+However, a tensor parallel group can extend across multiple servers.
+In such cases, ReCycle, along with Oobleck and Bamboo, classifies the entire group of tensor-parallel servers as a single unit of failure.
+4. ReCycle Design
 4.1. System Overview
-SlipStream consists of three key components, the Profiler, the Executors, and critically the Planner.
+Figure 8. ReCycle system overview.
+ReCycle consists of three key components, the Profiler, the Executors, and critically the Planner, as shown in Figure 8.
 Profiler.
-When a large training job is first submitted, SlipStream runs a short profiling job to collect key performance statistics such as the average micro-batch latency for the forward and backward passes, memory requirements for activations and gradients, and inter-node communication bandwidth.
+When a large training job is first submitted, ReCycle runs a short profiling job to collect key performance statistics such as the average micro-batch latency for the forward and backward passes, memory requirements for activations and gradients, and inter-node communication bandwidth.
 The profiling job executes a small number of training iterations, 100 by default, and typically takes a few minutes.
@@
 Planner.
-The Planner generates near-optimal schedules for the training job under various failure scenarios.
-It uses dynamic programming (DP) and mixed integer linear program (MILP) formulations that implement the SlipStream techniques (Section 3).
+The Planner generates pipeline schedules for the training job under various failure scenarios.
+It uses dynamic programming and mixed integer linear programming (MILP) to implement the ReCycle techniques (Section 3).
 We generate a pipeline schedule for each number of tolerated simultaneous failures.
 Each plan is agnostic to which specific worker(s) fail.
-It encodes the specific pipeline schedule including micro-batch assignment, the order of forward and backward tasks (e.g., ForwardPass, BackwardWeightPass), and the communication plan (e.g., RecvActiv, SendGrad, etc.).
+It encodes the specific pipeline schedule, specifying micro-batch assignments, the sequence of forward and backward tasks, and communication of activations and gradients.
 The plans are stored in distributed fault-tolerant storage (e.g., etcd) to be used by runtime Executors.
-By default, we generate up to $DP-1$ plans, which accounts for the $DP-1$ simultaneous failures that SlipStream can always handle.
-Since SlipStream can probabilistically sustain many more failures, we can also generate plans for up to a user-defined fault tolerance threshold.
-The Planner runs relatively quickly, requiring 53 minutes to generate plans for up to 512 concurrent failures for a cluster of 2048 GPUs.
-Nevertheless, the training job can begin training in parallel as most plans will not be needed immediately.
+By default, we generate up to $DP-1$ plans, which accounts for the $DP-1$ simultaneous failures that ReCycle can always handle.
+Since ReCycle can probabilistically sustain many more failures, we can also generate plans for up to a user-defined fault tolerance threshold.
 Executors.
-SlipStream operates with an Executor on each GPU node, tasked with implementing the training plan specific to that node.
+ReCycle operates with an Executor on each GPU node, tasked with implementing the training plan specific to that node.
 Overseeing these Executors, a centralized Coordinator orchestrates the overall training operation by distributing the appropriate execution plans to each Executor.
@@
 To enhance reliability, the Coordinator itself is safeguarded against failures through active replication.
-4.2. SlipStream Planner
-Given the cluster configuration, training job, profiling statistics, and set of failed workers, the Planner calculates a valid pipeline schedule that minimizes the training iteration latency using the fault-free workers.
-The Planner uses two phases, Normalization followed by Solution.
-SlipStream generates an execution plan for a given number of failures, regardless of which worker(s) fails, as solving ILP problems for the combinatorial number of possible failure locations is intractable.
-The first phase normalizes the specific failures by migrating failed nodes to a pre-determined and ideal set of locations within the pipeline schedule.
-For example, if any single node fails in the example in Figure 7, the Planner will always swap the node with an ideal location, e.g.
-$W_{2\_3}$.
-The Planner generates a mapping of the ideal migration locations for each possible number of simultaneous failures.
-Unlike current solutions which require a complete reconfiguration of a data-parallel pipeline, normalization only requires a point-to-point copy of model parameters to swap the location of two workers in the pipeline for each failure.
-The Solution Phase leverages an MILP to derive an optimized schedule for each normalized case.
-4.2.1. Normalization Phase
-The intuition behind the Normalization Phase is to a) distribute failures across different peer groups (to maximize fault tolerance and load balance extra work across peer groups) and b) move failures to later pipeline stages with more bubbles (to maximize the opportunity to hide overheads).
-Normalization relies on dynamic programming (DP) to determine the ideal pipeline location to migrate faults for given a number of failures.
-We use a DP algorithm as it can greedily determine the ideal migration strategy with very low runtime ($O(NF^{2})$ complexity for all solutions up to $F$ failures with $N$ stages per pipeline).
-The DP algorithm is based on a model that calculates the cost in terms of the extra bubbles used to accommodate re-routed micro-batches.
-DP Formulation
-Algorithm 1 Normalization Phase DP Algorithm
-1:$N\leftarrow$ Number of pipeline stages.
-2:$F\leftarrow$ Total number of failures.
-3:$C\leftarrow$ an Nx(F+1) array $\triangleright$ Costs
-4:$A\leftarrow$ an Nx(F+1) array $\triangleright$ Assignments
-5:for $i\in\{0,\dots,N-1\}$ do
-6:     for $f\in\{0,\dots,F\}$ do
-7:         if $i==0$ then
-8:              $C[i][f]=cost(i,f)$
-9:              $A[i][f]=[f]$
-10:         else
-11:              $C[i][f]=\min_{x\leq f}(C[i-1][f-x]+cost(i,x))$
-12:              $x=\operatorname*{arg\,min}_{x\leq f}(C[i-1][f-x]+cost(i,x))$
-13:              $A[i][f]=concat(A[i-1][f-x],x)$
-14:         end if
-15:     end for
-16:end for
-17:$R=A[N-1][F]$
-18:return $R$
-Given the total number of failures to handle, $F$, and the total number of pipeline stages within a data parallel pipeline, $N$, the DP algorithm returns a list $R$ of length $N$, such that $\Sigma_{i}R[i]=F$.
-Each index $R[i]$ specifies the ideal number of failures that is assigned to pipeline stage $i$ across all data parallel pipelines; the specific pipeline assignment does not affect performance and can be arbitrary.
-For example, if $R[3]=2$ for the example in Figure 7(a), the ideal migration would be to swap the failed nodes with any two nodes in stage $3$, for example $W_{0\_3}$ and $W_{2\_3}$.
-For each DP step, $C[i][f]$ represents the total cost, in terms of bubbles used, to handle $f$ failures at all stages between 00 and $i$.
-Meanwhile, each $A[i][f]$ contains a list that denotes the portion of the $f$ failures assigned to each stage from 00 to $i$, as described for $R$ – in other words $len(A[i][f])==i+1$.
-Each DP step uses a recurrence relation that aims to minimize the cost of handling $f$ failures within the first $i$ pipeline stages across all of the data parallel pipelines, while $A[i][f]$ denotes the assignment of failures that achieves this cost.
-The ideal assignment to handle $F$ failures across all workers is simply $R=A[N-1][F]$.
-To estimate the cost, $cost(i,x)$ is a function that returns the extra bubbles used to handle the extra micro-batches from $x$ failures in the $i$-th stage (i.e., peer group).
-While we can use the MILP described in Section 4.2.2 to estimate the cost, the Planner can also rely on a heuristic to estimate the cost in order to avoid longer solve times (as Algorithm 1 requires $O(NF)$ calls to the cost function).
-The heuristic captures the benefit of the techniques described in Section 3.
-It creates a pipeline schedule that is similar to 1F1B, except that the heuristic’s 1F1B backward pass only schedules $B_{input}$.
-Then, when there exists a gap (both in terms of time and memory) that is able to fit $B_{weight}$ (or the $F$ and $B_{input}$ stages are exhausted), $B_{weight}$ operations are scheduled opportunistically.
-4.2.2. Solution Phase
-Next, the Planner uses the normalized failure locations and the profiled statistics as inputs to an MILP which performs schedule optimization.
-At a high level, the MILP determines how to schedule re-routes micro-batches across peers of a failed worker alongside regular micro-batches in those peers (Adaptive Pipelining).
-The formulation accounts for the latency for communicating gradients across workers, task dependencies in forward and backward passes, and the memory usage in each worker.
-The definition of task dependencies also allows the solution to leverage Decoupled BackProp and Staggered Optimizer techniques.
-MILP Formulation
-We aim to minimize the makespan of a single training iteration which is repeated throughout the training process, while the constraints are task dependencies and the memory capacity of workers.
+4.2. Planner
+Given the cluster configuration, training job, profiling statistics, and failed workers, the Planner calculates an adaptive schedule that minimizes the training iteration latency using the fault-free workers.
+It does this in two phases, Failure Normalization followed by Adaptive Schedule Generation.
+To avoid solving an MILP for the combinatorial number of possible failure locations, the Failure Normalization phase first normalizes each given number of failures by calculating the suitable location within the pipeline schedule to migrate each failure.
+For example, if any single node fails in the example in Figure 7, the Planner will swap the node with the calculated ideal location, e.g., $W_{2\_3}$.
+Unlike current solutions which require a complete reconfiguration of a data-parallel pipeline (Jang et al., 2023), normalization only requires a point-to-point copy of model parameters to swap the location of two workers in the pipeline for each failure.
+The Adaptive Schedule Generation then leverages a MILP to derive an adaptive pipeline schedule for each normalized case.
+4.2.1. Failure Normalization
+Algorithm 1 Failure Normalization
+1:$DP\leftarrow$ Number of data-parallel pipelines.
+2:$PP\leftarrow$ Number of pipeline stages.
+3:$MB\leftarrow$ Number of microbatches per pipeline.
+4:$F\leftarrow$ Total number of failures.
+5:$O\leftarrow$ an PP $\times$ (F+1) array $\triangleright$ Rerouting Overheads
+6:$A\leftarrow$ an PP $\times$ (F+1) array $\triangleright$ Assignments
+7:
+8:procedure failure_reordering($DP,PP,MB,F$)
+9:     for $i\in\{0,\dots,PP-1\}$ do
+10:         for $f\in\{0,\dots,F\}$ do
+11:              if $i==0$ then
+12:                  $O[i][f]=\textsc{COST}(f)$
+13:                  $A[i][f]=[f]$
+14:              else
+15:                  $x=\operatorname*{arg\,min}_{x\leq f}\Bigl{(}O[i-1][f-x]$
+[…]
+17:                  $O[i][f]=O[i-1][f-x]+\textsc{COST}(x)$
+18:                  $A[i][f]=concat\Bigl{(}A[i-1][f-x],x\Bigr{)}$
+19:              end if
+20:         end for
+21:     end for
+22:return $A[PP-1][F]$
+23:end procedure
+24:
+25:procedure cost($f$)
+26:     if $f>0$ then
+27:         return $\min\Bigl{(}0,MB\times f\times 3-(DP-f)\times(PP-1)\times 3\Bigr{)}$
+28:     end if
+29:     return 00
+30:end procedure
+The intuition behind Failure Normalization, shown in Algorithm 1, is twofold: a) distribute failures across different peer groups to enhance fault tolerance and evenly balance the additional workload, and b) shift failures to later pipeline stages with more bubbles to maximize the opportunity to hide re-routing overheads.
+Failure Normalization uses dynamic programming to compute a migration strategy for handling a given number of failures.
+Given a total of $F$ failures and $PP$ pipeline stages, it returns a list $A$ of length $PP$, where the sum of elements in $A$ equals $F$.
+Each $A[i]$ specifies the number of failures assigned to pipeline stage $i$ across all data parallel pipelines, with the specific pipeline assignments being arbitrary and not impacting performance.
+For example, if $A[3]=2$, two failures would be assigned to stage 3, and we could swap the failed nodes with any two from that stage, such as $W_{0\_3}$ and $W_{2\_3}$ in Figure 7(b).
+The overhead of handling $f$ failures at stages from 00 to $i$ is represented by $O[i][f]$, while $A[i][f]$ indicates how the $f$ failures are distributed across these stages, with $A[i][f]$ being of length $i+1$.
+The recurrence relation minimizes overhead by finding the optimal failure assignment for $f$ failures across the first $i$ stages, and the best assignment for $F$ failures is given by $A[PP-1][F]$.
+To compute $COST(i,x)$, we estimate the additional time slots needed for handling extra micro-batches due to $x$ failures in the $i$-th stage using the heuristic from line 27.
+While the MILP in Section 4.2.2 could be used, Failure Normalization opts for the heuristic to reduce computation time.
+The complexity of determining the migration strategy for $F$ failures and $PP$ pipeline stages is $O(PP\times F)$.
+4.2.2. Adaptive Schedule Generation
+Next, the Planner uses normalized failure locations and profiled statistics as inputs for an MILP to generate adaptive schedule.
+At a high level, the MILP determines how to re-route micro-batches across peers of a failed worker, integrating both regular and re-routed micro-batches (Adaptive Pipelining).
+It considers communication latency $T_{comm}$, micro-batch latency for forward, backward input and weight pass – $T_{F}$, $T_{B_{Input}}$ and $T_{B_{Weight}}$ respectively, task dependencies in forward and backward passes, and memory usage, leveraging the Decoupled BackProp and Staggered Optimizer techniques.
 Notation.
-An operation within a training iteration can be explicitly identified across workers using a 5-tuple $(i,j,k,c,k_{s})$.
-$i$ represents the pipeline stage within each data parallel pipeline, while $k$ represents the operation’s original data parallel pipeline (prior to failures).
-$j$ is the micro-batch ID in the training iteration.
-$c$ identifies the different operations for each micro-batch, $c\in\{F,B_{input},B_{weight}\}$.
-Finally, $k_{s}$ denotes the peer pipeline in the event that a micro-batch originally intended for $k$ runs on $k_{s}$ (note that $k$ can equal $k_{s}$).
-For example, a micro-batch originally intended to execute on $W_{2\_3}$, but was re-routed to a peer $W_{1\_3}$ in Figure 7, would have $i=3$, $k=2$, and $k_{s}=1$.
-Thus, $(i,k_{s})$ uniquely identifies the worker for a given operation.
+Each operation in a training iteration is denoted by the 5-tuple $(i,j,k,c,k_{s})$.
+Here, $i$ represents the pipeline stage within each data parallel pipeline, and $k$ indicates the operation’s original data parallel pipeline before any failures.
+The variable $j$ denotes the micro-batch ID in the training iteration, and $c$ specifies the type of operation for each micro-batch, where $c\in\{F,B_{input},B_{weight}\}$.
+Finally, $k_{s}$ specifies the peer pipeline that executes a micro-batch originally intended for $k$; which can be the same as $k$.
+For instance, a micro-batch ID $14$ originally scheduled for $W_{2\_3}$ but rerouted to peer $W_{1\_3}$ would be identified with $i=3$, $j=14$, $k=2$, and $k_{s}=1$.
 Inputs.
-The MILP takes as input the micro-batch to worker assignments in an iteration.
-Specifically, the Planner takes the normalized failure locations and assigns all micro-batches corresponding to failed workers evenly across peer workers within the same peer group, alongside the original (non-peer) micro-batches.
-Thus, the Planner produces a binary mapping $S_{i,j,k}^{k_{s}}\in\{0,1\}$, which indicates if a micro-batch $(i,j,k)$ should be executed on pipeline $k_{s}$.
-Note that $S_{i,j,k}^{k}=1$ if the micro-batch executes on the original pipeline.
-We ensure that each operation is assigned only one pipeline, i.e., $\sum_{\begin{subarray}{c}k_{s}\end{subarray}}S_{i,j,k}^{k_{s}}=1$.
-The MILP also takes as input profiled statistics.
-First, $T_{c}$ represents the computational time for each operation $c$.
-$T_{comm}$ represents the communication latency to send activations or gradients across workers [^1] We use a single value as activations and gradients are the same size..
-Next, the MILP requires the change in memory utilization on worker $(i,k,k_{s})$ caused by executing an operation $(i,j,k,c,k_{s})$ as $\Delta M_{i,j,k,c}^{k_{s}}$.
+We use micro-batch assignments for each worker as inputs.
+The Planner reassigns micro-batches from failed workers to peer workers within the same group while retaining the original micro-batches.
+This results in a binary mapping $S_{i,j,k}^{k_{s}}\in\{0,1\}$, indicating whether a micro-batch $(i,j,k)$ should run on pipeline $k_{s}$.
+Each operation is assigned to exactly one pipeline: $\sum_{k_{s}}S_{i,j,k}^{k_{s}}=1$.
+Additionally, we use profiled statistics, which include $T_{c}$ for the computational time of each operation $c$, $T_{comm}$ for communication latency of activations or gradients [^1] We use a single value as activations and gradients are the same size., and $\Delta M_{i,j,k,c}^{k_{s}}$ for the change in memory utilization on worker $(i,k,k_{s})$ due to the execution of operation $(i,j,k,c,k_{s})$.
 $\Delta M_{i,j,k,c}^{k_{s}}=\begin{cases}A_{B}&\text{, if }c=F\text{ and }S_{i,j,k}^{k_{s}}=1\\ A_{B}-A_{B_{Input}}&\text{, if }c=B_{Input}\text{ and }S_{i,j,k}^{k_{s}}=1\\ -A_{B_{Weight}}&\text{, if }c=B_{Weight}\text{ and }S_{i,j,k}^{k_{s}}=1\\ 0&\text{, otherwise}\end{cases}$
-Here, $A_{B}$ is the profiled size of activation stored at the end of forward pass.
-$A_{B_{input}}$ and $A_{B_{Weight}}$ is size of the activation stored at the end of backward-input and backward-weight pass, respectively.
-Upon completion of backward-input pass and backward-weight pass, we free activation $A_{B_{Input}}$ and $A_{B_{Weight}}$ respectively.
+Here, $A_{B}$ is the profiled size of activation at the end of the forward pass.
+$A_{B_{input}}$ and $A_{B_{weight}}$ are the sizes of gradients at the end of the backward-input and backward-weight passes, respectively.
+We free $A_{B_{input}}$ and $A_{B_{weight}}$ after completing their respective backward passes.
 Variables.
-The goal of the MILP is to identify the exact ordering of operations, which can be used to generate a pipeline schedule.
-To do this, we introduce an ordering variable $O_{(i,j,k,c,k_{s})\rightarrow(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})}\in\{0,1\}$ between each pair of operations $(i,j,k,c,k_{s})$ and $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$.
-Each variable is 1 if $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled any time after $(i,j,k,c,k_{s})$ and 0 otherwise.
-Ordering is only required within a stage, and between computation phases of same micro-batch (e.g., between the $F$ and $B_{input}$ pass).
-Finally, we define the set of variables $E_{i,j,k,c}^{k_{s}}$ as the ending time of operation $(i,j,k,c,k_{s})$.
+We define a binary variable $O_{(i,j,k,c,k_{s})\rightarrow(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})}\in\{0,1\}$ to represent ordering between pair of operations $(i,j,k,c,k_{s})$ and $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$.
+This variable is 1 if $(i^{\prime},j^{\prime},k^{\prime},c^{\prime},k_{s}^{\prime})$ is scheduled after $(i,j,k,c,k_{s})$, and 0 otherwise.
+Ordering is needed only within a stage and between computation phases of the same micro-batch (e.g., $F$ and $B_{input}$).
+Additionally, $E_{i,j,k,c}^{k_{s}}$ represents the ending time of operation $(i,j,k,c,k_{s})$.
 Objective.
-The goal of the MILP is to minimize the makespan of a single training iteration (which is repeated throughout the training process).
-This can be done by determining the sequence and time of operations through the training pipeline via the sets of variables $O$ and $E$.
-This objective can thus be formulated as:
+Our objective is to minimize the makespan of a single training iteration while adhering to task dependencies and memory constraints.
+This involves determining the sequence and timing of operations throughout the training pipeline using the sets of variables $O$ and $E$.
+Thus, the objective can be formulated as follows:
 (1) $\min_{\begin{subarray}{c}O,E\end{subarray}}\qquad\max_{\begin{subarray}{c}i,j,k,k_{s}\end{subarray}}E_{i,j,k,B_{Weight}}^{k_{s}}$
@@
 […]
-Finally Equation 6 calculates the activation memory required at any given time on worker $(i,k_{s}^{\prime})$, and constrains operations such that the total memory is below an $M_{limit}$ to avoid OOMs.
-4.3. Implementation
-We implemented SlipStream on top of DeepSpeed (Rasley et al., 2020) and made the following additions to DeepSpeed to support SlipStream.
+Finally, Equation 6 calculates the activation memory required at any given time on worker $(i,k_{s}^{\prime})$, and constrains operations such that the total memory is below an $M_{limit}$.
+5. Implementation
+We implemented ReCycle on top of DeepSpeed (Rasley et al., 2020) and made the following additions to DeepSpeed to support ReCycle.
+Detecting failures.
+ReCycle uses existing hardware error detection mechanisms in GPUs (e.g.; SMBPBI APIs in NVIDIA GPUs) and software error detection in runtime systems (e.g.
+stderr logs in PyTorch).
+Workers send periodic heartbeats to a central driver, including worker information and hardware statistics from GPUs such as page retirement count, row-remapping stats, ECC correction stats, XID error logs, and stdout/stderr logs (Jiang et al., 2024).
+If the central driver detects abnormalities, it marks the worker as failed.
 Rerouting micro-batches to data parallel peers.
@@
 When a subsequent stage fails, ReRouteAct redistributes the micro-batches across the remaining peers in a round-robin fashion, while ReRouteGrad adjusts the gradient distribution to ensure that both the forward and backward processes of a micro-batch are handled by the same peer.
-This strategy maintains operational continuity in SlipStream under fault conditions, mirroring fault-free execution semantics.
+This strategy maintains operational continuity in ReCycle under fault conditions, mirroring fault-free execution semantics.
 These operators are integrated as pipeline instructions within the DeepSpeed execution engine.
 Decoupling Back Propagation in DeepSpeed.
-Our implementation of decoupled backpropagation within DeepSpeed centers around intercepting the weight gradient computations traditionally performed during the backpropagation phase.
-These computations are temporarily held in a newly created in-memory structure known as WeightGradStore, enabling SlipStream to defer weight gradient computation.
+Our implementation of Decoupled BackProp within DeepSpeed centers around intercepting the weight gradient computations traditionally performed during the backpropagation phase.
+These computations are temporarily held in a newly created in-memory structure known as WeightGradStore, enabling ReCycle to defer weight gradient computation.
 To facilitate this, two new pipeline instructions have been introduced into DeepSpeed’s execution engine: InputBackwardPass and WeightBackwardPass.
-These instructions are incorporated into SlipStream’s execution plan and are then processed by DeepSpeed’s execution scheduler, effectively managing the distinct phases of backpropagation.
+These instructions are incorporated into ReCycle’s execution plan and are then processed by DeepSpeed’s execution scheduler, effectively managing the distinct phases of backpropagation.
 Bypassing Optimizer Synchronizations.
-In DeepSpeed, each training iteration concludes with an all-reduce collective to synchronize gradients across data parallel peers, followed by a series of validation checks at each pipeline stage to ensure numerical stability, such as monitoring for gradient overflow.
-If any stage identifies an issue that could compromise numerical stability, the optimizer step for that iteration is omitted to avoid exacerbating the problem.
-Upon successful validation, the optimizer step is executed for each stage, and the training progresses to the next iteration in a synchronized manner across all stages.
-Table 1. Normalized training throughput with increasing failure rates. Fault-free training represents $1.00$, hence higher is better. Bamboo ran out of memory for GPT-3 3.35B and 6.7B.
+In DeepSpeed, each training iteration ends with an all-reduce collective to synchronize gradients across data-parallel peers, followed by validation checks at each pipeline stage to ensure numerical stability.
+If any stage detects a potential issue, the optimizer step for that iteration is skipped to prevent further problems.
+Upon successful validation, the optimizer step is executed, allowing synchronized progression to the next iteration.
+Table 1. Training throughput (samples/sec) with increasing failure frequency, higher is better. Bamboo ran out of memory for GPT-3 3.35B and 6.7B.
 Systems GPT-3 Medium GPT-3 3.35B GPT-3 6.7B
 Failure Frequency 6h 2h 30m 6h 2h 30m 6h 2h 30m
-Bamboo (Thorpe et al., 2023) 0.71 0.69 0.55 OOM OOM OOM OOM OOM OOM
-Oobleck (Jang et al., 2023) 0.99 0.92 0.71 0.98 0.90 0.66 0.93 0.87 0.52
-SlipStream 0.99 0.92 0.81 0.98 0.95 0.85 0.97 0.91 0.66
-In contrast, SlipStream employs a staggered timing approach for the optimizer steps across different pipeline stages to enhance scheduling efficiency.
+Fault-Free DeepSpeed (Rasley et al., 2020) 27.58 14.87 5.33
+Bamboo (Thorpe et al., 2023) 19.47 18.98 15.24 OOM OOM OOM OOM OOM OOM
+Oobleck (Jang et al., 2023) 27.26 25.37 19.47 14.55 13.44 9.78 4.98 4.65 2.78
+ReCycle 27.27 25.42 22.27 14.59 14.17 12.63 5.17 4.85 3.53
+In contrast, ReCycle uses a staggered timing approach for optimizer steps across different pipeline stages to improve scheduling efficiency.
 This method of staggering is not conducive to cross-stage synchronization for validating numerical stability prior to executing the optimizer.
-To accommodate staggered execution while maintaining validation integrity, numerical validations are shifted from a pre-step to a post-step process.
-Following the all-reduce collective, each stage immediately conducts its own local validation checks but does not wait for downstream stages to complete their validations.
-Each stage proceeds with its optimizer step based on its individual validation results and those of preceding stages.
-If a downstream stage fails its validation, a rollback is initiated across all stages before advancing to the next iteration.
-Notably, for many optimizers, including the widely-used AdamW (Loshchilov and Hutter, 2017), such rollbacks do not entail additional memory costs because the operations involved are arithmetically reversible.
-This adjustment allows SlipStream to retain the advantages of staggered operations while protecting the training process from potential instabilities.
+To address this, ReCycle shifts numerical validations from a pre-step to a post-step process.
+After the all-reduce collective, each stage performs its own local validation checks without waiting for downstream stages.
+Each stage executes its optimizer step based on its own validation results and those of preceding stages.
+If a downstream stage fails validation, a rollback occurs across all stages before moving to the next iteration.
+Notably, for many optimizers, including the commonly used AdamW (Loshchilov and Hutter, 2017), these rollbacks incur no additional memory costs due to the arithmetic reversibility of the operations.
+This adjustment enables ReCycle to leverage staggered operations while safeguarding the training process from potential instabilities.
 6. Evaluation
 Figure 9.
-Training throughput, normalized to the fault-free case, for the GPT-3 Medium and GPT-3 6.7B models over the GCP trace.
-In 8b and 8c, the dashed lines represent the average training throughput achieved by each system within the 6h period.
+Training throughput (samples/sec), higher is better, for the GPT-3 Medium and GPT-3 6.7B models over the GCP trace.
+In 9b and 9c, the dashed lines represent the average training throughput achieved by each system within the 6h period.
 6.1. Experimental Setup
 Cluster Setup.
-We performed real-world experiments on a cluster of 32 NVIDIA A100 80GB GPUs in Azure.
-Specifically, we used four Standard_NC96ads_A100_v4 instances, each with 8 GPUs, 96 vCPUs, and 880 GB of host memory.
-Each node has an 600 GB/s NVLink intra-node interconnect, as well as a 640 Gbps inter-node interconnect over 8 NICs.
-We performed scalability experiments using a simulator, which we present and validate in Section 5.3.
+We conducted real-world experiments on a 32-GPU cluster featuring NVIDIA A100 80GB GPUs in Azure, utilizing Standard_NC96ads_A100_v4 instances (8 GPUs, 96 vCPUs, and 880 GB memory each).
+Each node includes a 600 GB/s NVLink intra-node interconnect and a 640 Gbps inter-node interconnect across 8 NICs.
+Scalability experiments were conducted using a simulator, discussed in Section 6.3.
 Baselines.
-We compared SlipStream to the two state-of-the-art baselines discussed in Section 2.3 – Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023) – with all of their respective optimizations enabled.
-We report the fault-free throughput achieved by DeepSpeed (Rasley et al., 2020) using an 1F1B pipeline schedule (Narayanan et al., 2019).
-All baselines train on the same Azure cluster.
+We evaluated ReCycle against two state-of-the-art baselines discussed in Section 2.2.3 – Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023).
+Both baselines were tested with all their optimizations enabled.
+Additionally, we report the fault-free throughput achieved by DeepSpeed (Rasley et al., 2020) using a 1F1B pipeline schedule (Narayanan et al., 2019).
+All experiments were conducted on the same Azure cluster for consistency.
 Workloads.
-We evaluated all systems using the Megatron (Shoeybi et al., 2019) implementation of GPT-3 (Brown et al., 2020).
-We used three GPT-3 sizes: Medium (350M), 3.35B, and 6.7B.
-We used (PP, DP) degrees of (2, 16), (4, 8), and (8, 4) for the Medium, 3.35B, and 6.7B models, respectively.
-All models used a TP degree of 1.
-We trained on wikitext (Merity et al., 2016) using (batch, micro-batch) size of (8192, 8), (1024, 1), and (1024, 1) for the Medium, 3.35B, and 6.7B models, respectively.
-Unless otherwise noted, all real-world experiments ran for 6 hours.
+We evaluated all systems using the Megatron (Shoeybi et al., 2019) implementation of GPT-3 (Brown et al., 2020), with three model sizes: Medium (350M), 3.35B, and 6.7B.
+The applied (PP, DP) degrees were (2, 16), (4, 8), and (8, 4), respectively, with a TP degree of 1 for all models.
+Training was conducted on wikitext (Merity et al., 2016) with batch and micro-batch sizes of (8192, 8) for Medium, and (1024, 1) for both 3.35B and 6.7B models.
+All real-world experiments, unless stated otherwise, ran for 6 hours with 32 workers.
 6.2. Training Throughput Under Failures
-How well does SlipStream handle failures compared to baselines?
-For each 32-GPU training job on Azure, we instrumented a random GPU to fail at regular and increasing intervals: 6 hours, 2 hours, and 30 minutes, which is representative of real-world failure rates (Jeon et al., 2019; Weng et al., 2022), starting at the beginning of the training job ($t=0$).
-Note that we treat each individual GPU as a worker.
-Table 1 shows the average training throughput achieved by each system across models and failure rates normalized to the fault-free throughput achieved by DeepSpeed ($1.00$).
-Note that GPUs are not replaced after failure, hence the total number of workers monotonically declines throughout each experiment.
-With a failure every 30 minutes, the percentage of functional GPUs by the end of the 6 hour run is 62.5%.
-Hence, we expect the average throughput with a failure every 30min to be relatively modest compared to the fault-free case.
-Because Bamboo relies on redundant computation, it rapidly exhausts GPU memory with its additional copy of model states and activations.
-Hence, Bamboo fails to train the two larger models.
-Even for the smallest model, redundant computation introduces significant throughput degradation.
-A single GPU failure (3% of total GPUs) in the 6h job causes a 29% throughput drop.
-Bamboo actually penalizes fault-free execution by nearly the same percentage compared to DeepSpeed.
-Oobleck can handle all model sizes and achieves significant throughput improvements over Bamboo.
-However, as the frequency of failures and model size increase, Oobleck’s performance degrades.
-As failures increase, Oobleck’s heterogeneous pipelines become more imbalanced, degrading the throughput of the synchronous training job.
-As the model size increases, the latency of reconfiguring all of the stages in the pipeline with the failed GPU gets higher.
-SlipStream is able to efficiently continue training under failures, out-performing or matching the throughput of both baselines across all models and failure rates.
-Adaptive Pipelining, Decoupled BackProp, and Staggered Optimizer are effective in optimizing the use of compute and memory resources across peers of failed GPUs.
-Moreover, SlipStream recovers from a failure quickly as it migrates model parameters to only a single GPU during normalization (see Section 4.2.1).
-Under frequent failures (every 30m) and no repairs, SlipStream achieves 85% and 66% of the fault-free training throughput for GPT-3 3.35B and 6.7B respectively.
-SlipStream out-performs Oobleck and Bamboo by up to 1.29×1.29$\times$1.29 × and 1.46×1.46$\times$1.46 ×, respectively for these experiments.
-Note that, unlike Bamboo, both SlipStream and Oobleck introduce no penalty over DeepSpeed for fault-free execution.
-What is SlipStream’s throughput advantage in an realistic training scenario?
-Next, we evaluated a realistic training scenario where GPUs fail and re-join based on a real-world trace.
-We used a trace of node failures/repairs from GCP instances used by Bamboo (Thorpe et al., 2023).
-We observed similar results for Bamboo’s AWS trace, but omitted the plots due to space constraints.
-We replayed the events in the GCP trace over a 6h training run on our Azure cluster.
-At the time of the experiment, we had 24 GPUs available (instead of 32 GPUs).
-Figure 8 shows the number of available GPUs over time according to the trace (max 24, min 15).
-Figure 8 also plots the training throughput achieved by Bamboo, Oobleck, and SlipStream replaying the trace for a GPT-3 Medium and GPT-3 6.7B training job.
-Solid lines represent instantaneous training throughput, while dashed lines show the average training through for each system over the 6h period.
-Bamboo cannot train GPT-3 6.7B due to memory pressure.
-SlipStream achieves a 1.64×1.64$\times$1.64 × higher performance (average throughput) compared to Bamboo on GPT-3 Medium and a 1.46×1.46$\times$1.46 × higher performance than Oobleck on GPT-3 6.7B.
-As discussed above, Bamboo suffers from the overheads of redundant computations.
-Oobleck requires large stalls due to parameter re-shuffling both when GPUs fail and when they rejoin, causing frequent and large drops in throughput.
-Meanwhile, during periods of stability, Oobleck suffers from imbalanced heterogeneous pipelines as discussed above.
-SlipStream provides the highest and most consistent training throughput across all portions of the trace.
-5.3. SlipStream Scalability
+How well does ReCycle handle failures compared to baselines?
+We first evaluate the average training throughput of Bamboo, Oobleck and ReCycle on various failure scenarios.
+We set the frequency of failures from once every 6 hours to once every 30 minutes to cover a wide spectrum of environments (Jeon et al., 2019; Weng et al., 2022).
+We monotonically reduce the number of available workers without recovery, so the total number of workers steadily decreases over the course of each experiment – for example in the 30m case, only $62.5\%$ of the workers (20 out of 32 workers) remain at the end of training.
+Table 1 presents the average throughput for various failure frequencies and model sizes.
+Bamboo suffers from static overhead due to redundant computations and additional model state copies, which quickly depletes GPU memory.
+This leads to its inability to train larger models and a significant drop in throughput for even the smallest model—resulting in a 29% reduction in throughput in the 6-hour failure frequency scenario.
+In contrast, Oobleck effectively manages all model sizes and demonstrates throughput improvements over Bamboo.
+However, Oobleck’s performance declines with increasing failure frequency and model size due to imbalanced heterogeneous pipelines and higher reconfiguration latency.
+ReCycle effectively continues training despite failures, consistently matching or exceeding the throughput of both baselines across all models and failure rates.
+Unlike Bamboo, the techniques utilized in ReCycle—Adaptive Pipelining, Decoupled BackProp, and Staggered Optimizer—introduce no static overhead, enabling uninterrupted training for larger models and increasing throughput by up to 1.46×1.46$\times$1.46 × for GPT-3 Medium.
+Additionally, ReCycle has reduced reconfiguration overhead; during failure normalization, it requires parameter migration for at most one GPU, compared to the need for reconfiguring an entire data pipeline with Oobleck.
+This efficiency results in up to 1.29×1.29$\times$1.29 × higher training throughput.
+What is ReCycle’s throughput advantage in dynamic training scenarios?
+We next assessed ReCycle’s performance in a dynamic training scenario characterized by GPU failures and re-joins, using a real-world failure trace.
+This trace, derived from GCP instances utilized by Bamboo (Thorpe et al., 2023) and Oobleck (Jang et al., 2023), was replayed over a 6-hour training run on our Azure cluster.
+While we observed similar outcomes with Bamboo’s AWS trace, we opted to omit those plots due to space limitations.
+During the experiment, we had 24 GPUs available instead of the originally planned 32.
+Figure 9a illustrates the fluctuating number of available GPUs over time, ranging from a maximum of 24 to a minimum of 15.
+Unlike the previous experiment, this scenario frequently saw GPUs being removed and reintroduced into the cluster.
+In Figure 9b and 9c, we present the training throughput achieved by Bamboo, Oobleck, and ReCycle while replaying the trace for both GPT-3 Medium and GPT-3 6.7B training jobs.
+The solid lines indicate instantaneous training throughput, while the dashed lines represent the average training throughput for each system throughout the 6-hour period.
+Notably, Bamboo is unable to train GPT-3 6.7B due to memory constraints.
+ReCycle demonstrates a performance increase of 1.64×1.64$\times$1.64 × in average throughput compared to Bamboo on GPT-3 Medium, and a 1.46×1.46$\times$1.46 × improvement over Oobleck on GPT-3 6.7B.
+As mentioned earlier, Bamboo incurs overhead from redundant computations, while Oobleck experiences significant stalls due to parameter re-shuffling during GPU failures and re-joins, leading to substantial drops in throughput.
+Additionally, during stable periods, Oobleck grapples with imbalanced heterogeneous pipelines.
+In contrast, ReCycle consistently delivers the highest and most stable training throughput throughout the entire trace.
+6.3. ReCycle Scalability
 Table 2. Gap between real-world and simulated throughput across various models and failure rates.
@@
 GPT-3 6.7B +3.94% +2.71% -1.86% -0.85%
-Since we do not have access to a cluster of thousands of GPUs, we built a simulator that can calculate the training throughput of a model and cluster configuration, given an execution plan.
-The simulator uses real-world profiled statistics for each pipeline operation of the respective model.
-To demonstrate that the simulator yields accurate results, Table 2 shows the gap between the simulated throughput and measured real-world througpught on Azure across different failure rates for the three GPT-3 models.
-The maximum difference is $5.98\%$.
-These differences arise due to minor fluctuations in the execution time of NCCL collectives, which do not cause impactful discrepancies when evaluating SlipStream.
-How well does SlipStream scale to large clusters and models?
-Using the simulator, we explored how well SlipStream handles failures for large models training on larger clusters.
-Figure 9 shows of the simulated throughput achieved by SlipStream across four different GPT-3 models with up to 145.6B parameters, normalized to the simulated fault-free throughput achieved by DeepSpeed’s 1F1B schedule.
-We also show the fault-scaled throughput, which is obtained by multiplying the fault-free throughput by the percent of non-failed GPUs.
-Note that we report the steady-state throughput with 1%, 5%, and 10% failed GPUs, not the throughput as some number of GPUs fails and rejoins.
-We simulated a cluster of (# GPUs, # stages per pipeline, and # pipelines) = (256, 8, 32), (512, 16, 32), (1024, 32, 32) and (1536, 64, 24) for the 18.4B, 39.1B, 76.1B, and 145.6B model, respectively, which reflect the larger cluster sizes needed to train large models.
+Due to the unavailability of a cluster with thousands of GPUs, we developed a simulator capable of calculating the training throughput based on a model and cluster configuration, given a specific execution plan.
+The simulator leverages real-world profiled statistics for each pipeline operation associated with the respective model.
+To validate the simulator’s accuracy, Table 2 presents the differences between the simulated throughput and the measured real-world throughput on Azure across various failure rates for three GPT-3 models, with a maximum discrepancy of $5.98\%$.
+These variations are mainly due to minor fluctuations in the execution time of NCCL collectives, but they have little impact on ReCycle’s performance.
+How effectively does ReCycle scale to large clusters and models?
+Using the simulator, we report the training throughput for GPT models with sizes of 18.4B, 39.1B, 76.1B, and 145.6B across different clusters: (256 GPUs, 8 stages per pipeline, 32 pipelines), (512 GPUs, 16 stages per pipeline, 32 pipelines), (1024 GPUs, 32 stages per pipeline, 32 pipelines), and (1536 GPUs, 64 stages per pipeline, 24 pipelines).
+Figure 10 presents the throughput normalized to the fault-free throughput achieved by DeepSpeed’s 1F1B schedule.
+Additionally, we introduce the fault-scaled throughput, calculated by multiplying the fault-free throughput by the percentage of operational GPUs.
+Notably, we focus on steady-state throughput for scenarios with 1%, 5%, and 10% GPU failures, rather than the throughput during periods of GPU failures and re-joins.
 Figure 10.
-Simulated throughput of SlipStream as model size increases, normalized to the simulated fault-free 1F1B throughput.
+Simulated throughput of ReCycle as model size increases, normalized to the simulated fault-free 1F1B throughput.
 The Fault-Scaled throughput is the fault-free throughput scaled by the percent of non-failed GPUs.
-SlipStream is effective at achieving high throughput across failure rates and model sizes.
-For the case of 1% failures (higher reliability), SlipStream handles failures at or better than the fault-scaled throughput across all models and cluster sizes, often with zero degradation.
-The bubbles in peer workers are more than sufficient to efficiently execute the work for failed GPUs using SlipStream’s optimizations (see Figure 6).
-As the number of failed GPUs increases, SlipStream is still able to sustain high throughput.
-At $5\%$ GPU failures, SlipStream achieves comparable performance to the fault-scaled throughput.
-Even at a large $10\%$ failure rate (e.g., 154 failed GPUs for GPT-3 145.6B), SlipStream enables clusters to continue training while only observing between $0.5\%$ and $11.5\%$ degradation to the fault-scaled case.
-This suggests that SlipStream will be effective for training jobs in supercomputing-scale clusters with dynamic resource availability (Jiang et al., 2024; He et al., 2023; Zu et al., 2024).
-5.4. SlipStream Performance Breakdown
-Figure 10. The contribution of the three optimization techniques to SlipStream’s training throughput.
+ReCycle demonstrates a strong capability to maintain high throughput across varying failure rates and model sizes.
+For a failure rate of 1%, ReCycle manages failures at or better than the fault-scaled throughput for all models and cluster sizes, frequently exhibiting no degradation in performance.
+The bubbles in peer workers are more than adequate to efficiently handle the workload of failed GPUs utilizing ReCycle’s optimizations (see Figure 6).
+As the number of failed GPUs increases, ReCycle continues to sustain high throughput levels.
+At a 5% failure rate, ReCycle’s performance is comparable to the fault-scaled throughput.
+Even at a significant 10% failure rate (e.g., 154 failed GPUs for GPT-3 145.6B), ReCycle allows clusters to continue training while observing only between $0.5\%$ and $11.5\%$ degradation from the fault-scaled case.
+This indicates that ReCycle is well-suited for training tasks in supercomputing-scale clusters with dynamic resource availability (Jiang et al., 2024; He et al., 2023; Zu et al., 2024).
+6.4. ReCycle Performance Breakdown
+Figure 11. The contribution of the three optimization techniques to ReCycle’s training throughput.
 Figure 12.
-Memory utilization for SlipStream across pipeline stages for GPT-3 6.7B with 30m failure rates compared to DeepSpeed (failure-free).
+Memory utilization for ReCycle across pipeline stages for GPT-3 6.7B with 30m failure rates compared to DeepSpeed (failure-free).
 The dashed line shows the 80GB memory capacity of the A100 GPU.
-How does each of SlipStream’s techniques contribute to performance?
-To understand the benefit (and necessity) of each of the SlipStream techniques described in Section 3, we performed an ablation study by repeating the real-world experiment on the Azure cluster with the GPT-3 models and 30min failure frequency.
-We modified the MILP formulation in the Planner to progressively enable optimizations.
-Figure 10 shows the throughput achieved in each step of the study, normalized to the fault-free throughput.
-Adaptive Pipelining allows SlipStream to continue training in the face of failures, but suffers from a large throughput degradation due to overheads caused by extra work in the peers of failed workers.
-By adding Decoupled BackProp, SlipStream boosts the training throughput by 63% to 118% by using bubbles in the cool-down phase of the pipeline schedule to mask this overhead (see Figure 3(b)).
-The addition of the Staggered Optimizer further improves throughput by 7% to 11% up to $11\%$ as it allows the use of bubbles from the warm-up phase of the pipeline schedule.
-Can SlipStream efficiently exploit unused GPU memory?
-SlipStream makes the key insight that GPU memory capacity is not uniformly utilized throughout the pipelined execution.
-Later pipeline stages inherently use less GPU memory.
-Decoupled BackProp utilizes this free GPU memory as a buffer to defer $B_{weight}$ computations (see Figure 5).
-To explore this in more detail, we analyzed the real-world training of the GPT-3 6.7B model with a 30m failure interval as in Section 5.2.
-Figure 11 shows the peak GPU memory utilization across the 8 stages of a pipelined execution, compared to the memory utilization of the failure-free DeepSpeed baseline.
-DeepSpeed, and any system that uses a pipeline schedule similar to 1F1B, does not fully utilize GPU memory, especially in later pipeline stages (Narayanan et al., 2019).
-This is because GPU workers that execute earlier pipeline stages must store a larger number of intermediate results, as shown in Figure 3.
-Figure 11 shows how SlipStream exploits this opportunity and uniformly utilizes nearly all GPU memory to optimize pipelined execution.
-The MILP formulation is able to carefully model and manage the additional memory requirements within GPU memory constraints and avoid memory exhaustion.
-Table 3. SlipStream Planner latency (in seconds) to find optimized schedules for up to $25\%$ of failed GPUs.
-DP PP Degree
-Degree 2 4 8 16 32 64
-2 5.24 5.41 15.48 42.73 129.57 459.57
-4 7.65 8.62 24.48 59.75 190.54 719.4
-8 8.56 9.57 30.74 81.42 294.21 1038.2
-16 11.84 13.8 51.53 116.8 452.98 1721.45
-32 11.71 14.87 67.24 278.64 942.42 3153.07
-What is the overhead of SlipStream’s Planner?
-To understand how fast the Planner can generate optimized schedules for large clusters, Table 3 shows the time required to generate all necessary schedules for up to $25\%$ of node failures across various parallelism strategies.
-Recall that the user can configure a smaller or larger fault-tolerance threshold.
-The reported latency includes both Planner phases.
-We report the results using Gurobi MILP solver (Gurobi, 2024) running on an 96-core CPU.
-We can generate all solutions to support up to $512$ failures in the 2048-GPU case ($DP=32$, $PP=64$) within 3153 seconds (52.5 minutes).
-Given that the training time for large models on thousands of GPUs is measured in weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020), the Planner’s latency is insigificant.
+How does each of ReCycle’s techniques contribute to performance?
+To assess the benefits and necessity of each technique in ReCycle, as outlined in Section 3, we conducted an ablation study.
+This involved repeating the real-world experiment on the Azure cluster with GPT-3 models, using a failure frequency of 30 minutes.
+We adjusted the MILP formulation in the Planner to progressively enable each optimization.
+Figure 11 shows the throughput achieved at each stage of the study, normalized to the fault-free throughput.
+Adaptive Pipelining allows ReCycle to continue training despite failures, but it experiences significant throughput degradation due to the overhead from additional work required by peers of the failed workers.
+By introducing Decoupled BackProp, ReCycle boosts the training throughput by 63% to 118%, effectively utilizing bubbles during the cool-down phase of the pipeline schedule to mask this overhead (see Figure 3(b)).
+Further enhancements come from the Staggered Optimizer, which boosts throughput by an additional $7\%$ to $11\%$ by leveraging bubbles from the warm-up phase of the pipeline schedule.
+Can ReCycle efficiently exploit unused GPU memory in hybrid parallelism?
+We recognize that GPU memory utilization varies across different stages in hybrid parallelism, with later pipeline stages typically requiring less memory.
+To leverage this, the Decoupled BackProp technique utilizes the available surplus GPU memory as a buffer, allowing for the postponement of $B_{weight}$ computations (see Figure 5).
+Figure 12 shows the peak GPU memory utilization across the eight stages of pipelined execution for ReCycle compared to the fault-free DeepSpeed, based on training the GPT-3 6.7B model with a 30-minute failure interval.
+DeepSpeed – and any system utilizing a pipeline schedule akin to 1F1B – fails to fully utilize GPU memory, especially in the later pipeline stages (Narayanan et al., 2019).
+This surplus arises because GPU workers in earlier stages must store more intermediate results, as shown in Figure 3.
+In contrast, ReCycle effectively capitalizes on this opportunity, achieving near-complete utilization of GPU memory to optimize adaptive pipelined execution in the presence of failures.
+The Planner’s MILP formulation is tailored to model and manage these additional memory requirements while adhering to GPU memory constraints, thereby preventing memory exhaustion.
+Figure 13. ReCycle Planner latency (in seconds) to find optimized schedules for up to $25\%$ of failed GPUs.
+What is the overhead of ReCycle’s Planner?
+To evaluate the latency of the Planner in generating adaptive schedules for large clusters, Figure 13 shows the time required to generate all necessary schedules for up to 25% node failures across various hybrid parallelism strategies.
+Users have the flexibility to configure the fault-tolerance threshold to suit their needs, whether smaller or larger.
+The reported latency encompasses both phases of the Planner and was measured using the Gurobi MILP solver (Gurobi, 2024) on a 96-core CPU.
+In a training setup with 2048 GPUs (with $DP=32$ and $PP=64$), the Planner can generate all adaptive schedules to accommodate up to $512$ failures in just $3153$ seconds (or $52.5$ minutes).
+Considering that training large models on thousands of GPUs typically takes weeks (Smith et al., 2022; Jiang et al., 2024; Naumov et al., 2020), the Planner’s latency is negligible, accounting for less than $0.1\%$ of the overall training time.
 7. Related Work
@@
 Alternatively, model parallelism (Krizhevsky, 2014; Shoeybi et al., 2019; Shazeer et al., 2018) involves distributing the components of a deep neural network (DNN) model across multiple GPU devices, allowing each device to handle a specific subset of the model’s parameters for all input data.
-Recently, pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023) has emerged as a technique for training large models by dividing the model’s layers among different workers and utilizing microbatches to efficiently use the available computational resources.
-Prominent deep learning training frameworks like PyTorch (PyTorch, 2024), DeepSpeed (Rasley et al., 2020), and Megatron (Narayanan et al., 2021) have adopted hybrid-parallelism, an approach that integrates data parallelism, model parallelism, and pipeline parallelism.
+Recently, pipeline parallelism (Huang et al., 2019; Narayanan et al., 2019; Li and Hoefler, 2021; Liu et al., 2023; Kim et al., 2023) has emerged as a technique for training large models by dividing the model’s layers among different workers and utilizing micro-batches to efficiently use the available computational resources.
+Prominent deep learning training frameworks like PyTorch (pyt, 2024), DeepSpeed (Rasley et al., 2020), and Megatron (Narayanan et al., 2021) have adopted hybrid-parallelism, an approach that integrates data parallelism, model parallelism, and pipeline parallelism.
 This integration facilitates training on a massive scale while enhancing computational and memory efficiency.
@@
 Alpa (Zheng et al., 2022) automatically optimizes inter- and intra-operator parallelism using a hierarchical ILP formulation.
-Tofu (Wang et al., 2019a) uses DP to optimally partition tensor operations in a single node.
+Tofu (Wang et al., 2019a) employs dynamic programming to optimally partition tensor operations in a single node.
 FlexFlow (Jia et al., 2018) uses a randomized search algorithm to quickly find parallelism strategies.
-TensorOpt (Cai et al., 2022) introduces a DP that can optimize parallelism for multiple resource dimensions (e.g., memory and compute).
-Piper (Tarnawski et al., 2021) proposes a two-level DP algorithm to find optimal hybrid parallelism strategies.
-SlipStream uses a DP algorithm and an MILP which exploit key functional redundancies in hybrid parallelism to explicitly improve training throughput under failures.
+TensorOpt (Cai et al., 2022) introduces a dynamic programming approach capable of optimizing parallelism across multiple resource dimensions, including memory and compute.
+Piper (Tarnawski et al., 2021) proposes a two-level dynamic programming algorithm to find optimal hybrid parallelism strategies.
+In contrast, ReCycle leverages a dynamic programming algorithm and mixed-integer linear programming (MILP) to exploit key functional redundancies in hybrid parallelism, explicitly enhancing training throughput in the presence of failures.
 Elastic Training.
 Elastic training systems can dynamically adjust the resources allocated to a training job.
-Both Horovod (Sergeev and Del Balso, 2018) and Torch Distributed (Li et al., 2020) provide mechanisms to modify the number of workers, but require either a restart from a checkpoint or an expensive re-shuffling of model parameters.
+Both Horovod (Sergeev and Del Balso, 2018) and Torch Distributed (Li et al., 2020) offer mechanisms to modify the number of workers, but require either a restart from a checkpoint or an expensive re-shuffling of model parameters.
 CoDDL (Hwang et al., 2021), Optimus (Peng et al., 2018), OASiS (Bao et al., 2018), and Themis (Mahajan et al., 2020) are ML cluster schedulers that can dynamically allocate resources across multiple DNN training jobs.
@@
 (Or et al., 2020) auto-scale the number of workers for a training job.
-Varuna (Athlur et al., 2022) provides elastic training by leveraging spot instances, but requires restarts from checkpoints to handle unexpected pre-emptions.
-While some elastic training systems can affect model consistency by changing important hyperparameters such as batch size and learning rate, SlipStream ensures that operations are mathematically consistent regardless of the number of failures.
+Varuna (Athlur et al., 2022) provides elastic training by leveraging spot instances, but requires restarts from checkpoints to handle unexpected preemptions.
+While some elastic training systems can affect model consistency by changing critical hyperparameters such as batch size and learning rate, ReCycle guarantees mathematical consistency of operations, regardless of the number of failures.
 Fault-tolerant Training.
-DNN training systems widely leverage checkpoints as the basic mechanism for fault recovery (Rasley et al., 2020; Zu et al., 2024; Le Scao et al., 2022; Jiang et al., 2024; He et al., 2023; Beaumont, 2022; Jeon et al., 2019; Wang et al., 2019b).
-While naive checkpointing can lead to training stalls, recent work aims to reduce these overheads.
-CheckFreq (Mohan et al., 2021) dynamically adapts the checkpointing frequency to reduce overheads.
-Check-N-Run (Eisenman et al., 2022) quantizes in-frequently updated embedding table checkpoints for recommendation models.
-Gemini (Wang et al., 2023) introduces checkpoint traffic scheduling across the storage hierarchy (local CPU memory, remote CPU memory, and remote persistent storage).
+DNN training systems commonly use checkpoints for fault recovery (Rasley et al., 2020; Zu et al., 2024; Le Scao et al., 2022; Jiang et al., 2024; He et al., 2023; Beaumont, 2022; Jeon et al., 2019; Wang et al., 2019b), though naive checkpointing can cause stalls.
+Recent works like CheckFreq (Mohan et al., 2021), Check-N-Run (Eisenman et al., 2022), and Gemini (Wang et al., 2023) reduce overheads by adapting checkpoint frequency, quantizing embedding tables, and scheduling checkpoint traffic across the storage hierarchy, respectively.
 Megascale (Jiang et al., 2024) alleviates the storage bottleneck during recovery by sharing data between corresponding GPU workers across data parallel groups.
@@
 Bamboo (Thorpe et al., 2023) introduces redundant computation (RC) to provide resilience in the presence of frequent preemptions of training with spot instances.
-Oobleck (Jang et al., 2023) provides resiliency through creation of heterogenous pipelines.
-SlipStream matches or outperforms them for a wide range of model sizes and failure frequencies.
+Oobleck (Jang et al., 2023) provides resiliency through creation of heterogeneous pipelines.
+ReCycle matches or outperforms them for a wide range of model sizes and failure frequencies.
 8. Conclusion
-SlipStream enables efficient DNN training in the presence of failures without the use of spare resources.
-SlipStream exploits the functional redundancies inherent to hybrid parallel training systems used for large DNNs.
-It introduces novel techniques to minimize the training throughput degradation of sustaining failures by exploiting unused resources (e.g., pipeline bubbles).
-We evaluated an end-to-end SlipStream prototype and showed it can tolerate a high numbers of concurrent failures for systems and models of various sizes.
-SlipStream improves training throughput under failures over Oobleck and Bamboo by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 × respectively.
+ReCycle enables efficient DNN training in the presence of failures without relying on spare resources.
+By leveraging the functional redundancies inherent in hybrid parallel training systems for large DNNs, ReCycle introduces innovative techniques to minimize the degradation in training throughput caused by failures.
+It does this by utilizing unused resources, such as pipeline bubbles.
+We evaluated an end-to-end prototype of ReCycle and demonstrated its ability to tolerate a high number of concurrent failures across systems and models of varying sizes.
+Compared to Oobleck and Bamboo, ReCycle improves training throughput under failure conditions by up to 1.46×1.46$\times$1.46 × and 1.64×1.64$\times$1.64 × respectively.
+Acknowledgements
+We are grateful to the anonymous reviewers and to our shepherd, Laurent Bindschaedler, whose comments have greatly helped improve this paper.
+This research was partly supported by the Stanford Platform Lab and its affiliates, and by ACE, one of the seven centers in JUMP 2.0, a Semiconductor Research Corporation (SRC) program sponsored by DARPA.
+Athinagoras Skiadopoulos was supported by a Stanford Graduate Fellowship.
+Mark Zhao was supported by a Stanford Graduate Fellowship and a Meta PhD Fellowship.